### v2

# Rethinking the Integration of Prediction and Planning in Deep Learning-Based Automated Driving Systems: A Review

Abstract
Automated driving has the potential to revolutionize personal, public, and freight mobility. Beside accurately perceiving the environment, automated vehicles must plan a safe, comfortable, and efficient motion trajectory. To promote safety and progress, many works rely on modules that predict the future motion of surrounding traffic. Modular automated driving systems commonly handle prediction and planning as sequential, separate tasks. While this accounts for the influence of surrounding traffic on the ego vehicle, it fails to anticipate the reactions of traffic participants to the ego vehicle’s behavior. Recent models increasingly integrate prediction and planning in a joint or interdependent step to model bi-directional interactions. To date, a comprehensive overview of different integration principles is lacking. We systematically review state-of-the-art deep learning-based prediction and planning, and focus on integrated prediction and planning models. Different facets of the integration ranging from model architecture and model design to behavioral aspects are considered and related to each other. Moreover, we discuss the implications, strengths, and limitations of different integration principles. By pointing out research gaps, describing relevant future challenges, and highlighting trends in the research field, we identify promising directions for future research.
Index Terms:
Automated Driving, Motion Prediction, Motion Planning, Deep Learning
{textblock*}
6.6in(0.95in,0.15in) SUBMITTED TO REVIEW AND POSSIBLE PUBLICATION. COPYRIGHT WILL BE TRANSFERRED WITHOUT NOTICE.
Personal use of this material is permitted. Permission must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
1 Introduction
Automated Driving (AD) remains a challenging endeavour. It is usually split into the subtasks of perception, prediction, planning, and control [1, 2, 3]. Perception processes sensor inputs to create a model of the environment. Prediction and planning build upon this model and make future motion forecasts for surrounding traffic agents and a plan for the controlled ego vehicle. Traditional, modular systems (cf. Fig. 1) address prediction and planning as separate tasks. The predicted behavior of surrounding traffic agents is leveraged to plan a suitable behavior for the ego vehicle. However, this sequential ordering is inherently reactive and cannot represent bidirectional interaction between the ego vehicle and other traffic agents [4]. In fact, prediction and planning are no sequential problems and should be tightly coupled in automated driving systems [5, 6, 7].
1.1 Scope
In this work, we review methods to integrate prediction and planning into an automated driving system (ADS).
Prediction is the task of anticipating intents and future trajectories of observed traffic agents [8], and planning is about finding the best possible trajectory w.r.t. previously defined criteria for a controlled vehicle [9].
Deep Learning (DL)-based methods follow a data-centric approach to tackle these problems [10] and have led to significant improvements in many fields [11]. Works in prediction and planning increasingly employ DL-based approaches as well [12, 13, 14, 15] and represent an alternative to traditional, rule-based methods [16, 17] and constrained optimization [18]. In this work, we focus on DL-based methods.
The architecture of an ADS can take different forms, which are shown in Fig 1. Black-box methods employ a single neural network for perception, prediction and planning (c.f. Fig 1(a)), so that depending on the architecture, no boundaries between the modules can be drawn [9, 19, 20]. In contrast, modular systems consist of clearly defined components for different tasks [21, 22, 23, 24, 25]. Traditionally, they are arranged in a sequential order (c.f. Fig 1(b)). In this work, we emphasize the shortcomings of this approach and highlight works that integrate prediction and planning in a more sophisticated way (c.f. Fig 1(c)) to enable bi-directional interaction of self-driving vehicles and surrounding traffic. We refer to the component of an automated driving system that integrates prediction and planning as an Integrated Prediction and Planning System (IPPS).
We limit the scope of our work to mixed traffic without Vehicle-to-X communication between traffic agents. Furthermore, we exclude pedestrian motion forecasting, which poses a different problem that is subject to weaker dynamic constraints compared to vehicles [26] and has been thoroughly surveyed [27, 28, 29].
1.2 Contribution and Structure
In a thorough approach, this survey first reviews DL-based prediction and planning before focusing on how to integrate these two tasks. While surveys on classical methods [30, 31, 14, 32], stand-alone DL-based prediction [33, 34, 35, 15] or planning [36], and end-to-end (E2E) AD [37] already exist, to the best of our knowledge we are the first to review the integration of DL-based prediction and planning. We summarize our contributions as follows:
  • We provide a comprehensive overview of common representations, architectures and state-of-the-art methods for DL-based prediction (Sec. 3) and planning (Sec. 4).
  • We propose a categorization for the integration of prediction and planning, which addresses model architecture, system design and behavioral aspects. Further, we investigate and analyze the interrelations between these categories and their impact on safety and robustness (Sec. 5).
  • We reveal gaps in state-of-the-art research and point out promising directions of future research based on the identified categorization (Sec. 6).
(a) Black box ADS
(b) ADS with sequential IPPS
(c) ADS with bidirectional IPPS
Figure 1: System architectures of automated driving systems. Dashed lines indicate, that due to the black-box nature of some E2E ADS, it can be unclear which network components do prediction and planning, so that no boundary around the IPPS can be drawn.
2 Task Definitions
In the following, we start by introducing the terminology and notation for the task definitions used in this work. We adopt a similar terminology to that proposed by [33] and partition the actors in a traffic scenario into the self-driving Ego Vehicle (EV) and the Surrounding Vehicles (SVs). The EV is equipped with sensors that provide information on the environment. The state history of vehicle $i$ over the time interval $t-t_{\textrm{obs}},\ldots,t$ is
$X_{i}=\{x_{t-t_{\textrm{obs}}},\ldots,x_{t-1},x_{t}\}.$ (1)
Each state $x$ comprises 2D or 3D positional information and further optional information like heading angle, speed, static attributes, or goal information in the case of the EV. Hence, $X_{\textrm{EV}}$ denotes the EV’s past states. Similarly,
$\overline{\textbf{X}}_{\textrm{SV}}=\{X_{1},X_{2},\ldots,X_{m}\}$ (2)
refers to the past states of all SVs. Analogously, the future states of vehicle $i$ within a prediction horizon of $t_{\textrm{pred}}$ are
$Y_{i}=\{y_{t+1},y_{t+2},\ldots,y_{t+t_{\textrm{pred}}}\}.$ (3)
and the future states of all SVs are $\overline{\textbf{Y}}_{\textrm{SV}}$. In the following, state sequences are also referred to as trajectories.
Additional scene information, such as a semantic map or traffic signs and traffic light states are represented by $I$.
Prediction. Following [8], we state trajectory forecasting as the task of estimating a probability distribution
$P_{\textrm{pred}}=P(\overline{\textbf{Y}}|\overline{\textbf{X}},I)$ (4)
that maps state histories of $m$ observed vehicles in $\overline{\textbf{X}}$ to the future trajectories of $n$ predicted vehicles in $\overline{\textbf{Y}}$. The distribution $P_{\textrm{pred}}$ accounts for inherent uncertainties in the forecasting task and is often modeled by a discrete set of samples with corresponding probabilities [38, 39, 40, 41]. Some methods omit scene information $I$ and infer trajectories based on a state history alone [42, 43, 44].
TABLE I: Forms of Prediction
$\overline{\textbf{Y}}$ $\overline{\textbf{X}}$ Task
$Y_{\textrm{SV}}$ $\overline{\textbf{X}}_{\textrm{SV}}$, $X_{\textrm{EV}}$ Single-Agent Prediction
$\overline{\textbf{Y}}_{\textrm{CV}}$ $\overline{\textbf{X}}_{\textrm{SV}}$, $X_{\textrm{EV}}$ Clique Prediction
$\overline{\textbf{Y}}_{\textrm{SV}}$ $\overline{\textbf{X}}_{\textrm{SV}}$, $X_{\textrm{EV}}$ Joint Prediction
[…]
Planning. The task is to find a single suitable trajectory for the EV that can be passed on to a downstream motion controller. Hence, we define planning to be a function $f$ that maps the observational inputs $\overline{\textbf{X}}_{\textrm{SV}}$, and $X_{\textrm{EV}}$ as well as the context information $I$ to a future trajectory $Y_{\textrm{EV}}$, i.e.
$Y_{\textrm{EV}}=f(X_{\textrm{EV}},\overline{\textbf{X}}_{\textrm{SV}},I).$ (5)
In many cases, the $f$ also uses the prediction $\overline{\textbf{Y}}_{\textrm{SV}}$, i.e.
$Y_{\textrm{EV}}=f(X_{\textrm{EV}},\overline{\textbf{X}}_{\textrm{SV}},I,\overline{\textbf{Y}}_{\textrm{SV}}).$ (6)
The definitions show that planning can be considered a special case of single-agent prediction where the output distribution models only a single trajectory. However, in contrast to prediction, the planning trajectory must be conditioned on a navigation goal. Moreover, it has to be stable in closed-loop, i.e. it must be kinematically feasible and result in safe and efficient driving behavior when fed to a downstream controller.
3 Prediction
The core of prediction is understanding how the driving scene will evolve. Sec. 3.1 discusses how different scene representations allow local and global interactions. Sec. 3.2 reviews which NN designs are used to model interactions and extract descriptive features. Finally, Sec. 3.3 shows how the extracted features are mapped to trajectory predictions and how multimodality is modeled.
3.1 Scene Representation
Figure 2: Comparison of rasterized and sparse representation of the same traffic scene. While the former represents information with fixed-size grids, the latter uses a set of vectors.
Representing a scene means extracting the relevant subset of all available information and transferring it into a format that subsequent processing steps can utilize. In the context of trajectory prediction for automated vehicles, agent states $\overline{\textbf{X}}$ and map I are the most important information to represent. Two major representations exist in DL-based approaches: Rasterized and sparse (cf. Fig. 2).
Rasterized. Rasterized representations use dense, fixed-resolution grid structures that often have multiple channels [49, 50, 48, 51, 52]. Each channel encodes different information on agent states. They are commonly used in combination with rasterized high-definition maps (HD maps). Typical modalities are occupancy maps [51], object detections [53], lane markings [54], motion history [55], semantic information [56], traffic signs [57], traffic light states [57], and route information [58] (c.f. Fig 2.
DESIRE [8] is an early work that uses a rasterized representation for DL-based prediction. A bird’s eye view (BEV) representation is employed to fuse multiple EV sensor inputs into a shared representation. This allows the combination of different sensing modalities and establishes a common coordinate system for all vehicles, facilitating to model interactions within the scene [59].
Additionally, dense grids are well-suited for processing with powerful Convolutional Neural Network (CNN) [60] architectures [58, 61, 62, 10, 63, 64, 65, 66, 67, 68, 69, 70, 71]. However, embedding all observations in a grid structure leads to information loss due to quantization errors [72]. Moreover, as discussed in Sec. 3.2, the locally restricted receptive field of CNNs can hamper interaction modeling.
Sparse representations use vectors to describe all objects in a scene. The per-object vectors are then either encoded jointly, e.g., with a transformer, or individually and then aggregated, e.g., with Graph Neural Networks (GNN) [73]. Objects are represented by polygons, or point sets [6, 74, 41, 75, 61, 76, 38, 77]. The map is commonly described by a set of lanes represented by polylines and an adjacency matrix. Objects and map elements are then encoded into fixed-size latent features, for instance, via Multi-Layer Perceptrons (MLP) [78] or Recurrent Neural Networks (RNN) [79, 39]. To encode sequential signals like state histories $\overline{\textbf{X}}$, 1D CNNs [80], RNNs like LSTMs [81] and GRUs [82], and Transformers [83] are applied [41, 84, 76].
The timeline in Fig. 3 chronologically depicts impactful prediction methods and exhibits a shift towards sparse representations during the past years. Diehl et al. [72] proposed a graph representation for vehicle interaction modeling. Based on their work, two approaches emerged: Homogeneous and heterogeneous graphs. Homogeneous graphs encode objects of one class, for example, the road network [74, 41], whereas heterogeneous graphs combine different object classes in one graph [40, 85, 86, 87]. However, their work is focused on highway data and does not include map information. In contrast, VectorNet [74] proposes to represent all map elements and agents by a set of points. Agents are represented by a set of points along their motion history, while lanes are described by polyline points. Small objects, such as traffic signs, are represented by a single point, and extended elements are given by polygons defined by multiple points. Vectornet then processes each object with a homogeneous subgraph called polyline subgraph. A per-object feature is obtained by max-pooling. Next, all object features are aggregated in a heterogeneous global graph. Finally, the TV trajectory is decoded from the encoding of the node corresponding to the TV. Trajectron++ [40] and MFP [88] combine rasterized representations for HD maps with sparse representations for objects. Further, a few works use voxel representations that differ in their degree of sparsity [89][90].
MultiPath++ [38] is the successor of MultiPath [56] and switches from rasterized to sparse representations. Their results show that the shift from rasterized to sparse can increase performance. The overall development in the field confirms this trend (c.f. Fig. 3).
Coordinate frame. Irrespective of the scene representation itself, the coordinate frame is a fundamental design choice.
Applying a global coordinate system with a fixed viewpoint for the whole scene [91, 46, 92, 6] is computationally efficient. However, the frame is not viewpoint-invariant, i.e., the predictions change if the coordinate system is moved within the scene. This leads to low sample efficiency and impairs generalization [93]. Even fixing the coordinate origin to the EV center does not lead to viewpoint-invariant predictions when predicting other agents than the EV [94].
In contrast, other models [95, 96, 74, 54] process a scene from the viewpoint of each agent. Thus, the processing is viewpoint-invariant. On the downside, computational complexity scales linearly with the number of agents and quadratically with the number of interactions [6].
An alternative way to achieve viewpoint invariance are pairwise relative coordinate systems introduced in GoRela [97], which describe the relation between agents instead of using a fixed coordinate frame. This allows the computation of the relation between static objects offline and greatly reduces computation. PEP [94] further reduces computation by using center-relative instead of pairwise relative coordinates, i.e. shifting the coordinate frame into the mean position of all processed agents before each transformation.
Another choice to increase robustness are Frenet representations that decompose positions into progress along and lateral offset from a lane [98, 99]. Either a single Frenet frame defined by the target vehicle is used for all agents [100] or each agent is represented in an individual coordinate frame [84, 101].
2016201720182019202020212022202320241988DESIRE [8]Fast and Furious $\uparrow$ [102]R2P2 $\downarrow$ [103]MTP [54]IntentNet [49]MATF [89]MultiPath $\uparrow$ [56]MFP [88]CoverNet $\downarrow$ [55]Trajectron++ [40]VectorNet [74]LaneGCN $\uparrow$ [41]TNT $\downarrow$ [75]mmTransformer [76]HOME $\uparrow$ [104]AgentFormer $\downarrow$ [105]DenseTNT $\uparrow$ [106]GOHOME $\downarrow$ [84]MultiPath++ [38]SceneTransformer $\uparrow$ [6]HEAT $\downarrow$ [107]HiVT [108]Wayformer [109]Motion Transformer [110]GoRela [97]LaFormer [111]MotionDiffuser [112]Trajeglish $\uparrow$ [113]BAT $\downarrow$ [114]Scene Representation (3.1)RasterizedSparseMixedInteraction Model (3.2)RNNCNNGNN / AttentionTransformerTrajectory Decoding (3.3)Trajectory RegressionAnchor Trajectory ClassificationEndpoint and CompletionInitial Trajectory and RefineAuto-Regressive FormulationPrediction2016201720182019202020212022202320241988ALVINN [9]Deepdriving [115]End-to-End-CNN [19]FCN-LSTM [116]CIL [117]Camera360 [118]Seg2WP [22]CAL [119]DeepImitative [62]DOCP [10]ChauffeurNet [58]DILP [63]CILRS [64]NMP [61]UD-CIL $\uparrow$ [23]NEAT $\uparrow$ [120]JLTP $\downarrow$ [121]LBC [67]LSD [122]PILOT [123]Transfuser [20]UrbanDriver $\downarrow$ [24]LAV [124]PlanT [125]PDM-Open [126]Input Space (4.1)LatentInterpretableOutput Space (4.2)ActionTrajectoryGoal Conditioning (4.3)NoneSubmodulesInput FeatureRoute CostRoute AttentionPlanning Paradigm (4.4)RegressionCost FunctionHybridPlanning2016201720182019202020212022202320241988PAC [127]HRI [128]PRECOG [65]PiP [129]DSDNet $\uparrow$ [68]P3 $\downarrow$ [66]LookOut $\uparrow$ [69]MP3 $\downarrow$ [70]SafetyNet [130]DIPP [7]SafePathNet [131]GameFormer $\uparrow$ [132]PDM-Hybrid $\uparrow$ [126]UniAD $\downarrow$ [133]TPP [134]VAD $\downarrow$ [135]DTPP [136]Ego-Agents Relation (5.2)UnknownRobot LeaderHuman LeaderJoint PlanningCo-LeaderSafety & Contingency (5.3)No CostMarginalizedWorst-CaseJoint CostContingentGoal Conditioning (4.3)NoneSubmodulesInput FeatureRoute CostRoute AttentionPlanning Paradigm (4.4)RegressionCost FunctionHybridManual IPPS
Figure 3: Contributions that got attention in the research fields of DL-based prediction, planning, and manual IPPS for automated driving. Nodes, stem, and label of each entry encode different aspects that are explained below the timelines. Works marked with $\uparrow$ and $\downarrow$ were slightly moved up and down the timeline, respectively, to avoid overlaps.
3.2 Interaction Modeling
A decisive aspect of prediction is to model the interaction between traffic agents represented in the scene [4, 137]. This includes interaction with map elements and other traffic agents. Modeling the relation of vehicles to static map elements like lane markings, traffic signs, etc. is important to understand feasible corridors in which trajectories could be localized. Interaction with other traffic agents are crucial to understanding social interaction.
RNNs. A few early prediction models like DESIRE [8] or Trajectron++ [40] combine RNNs with an aggregation operator like spatial pooling or attention. They either extract temporal per-agent features and then aggregate them across all agents or aggregate information first and extract joint temporal features afterwards.
CNNs. Alternatively, Fast and Furious [102], MTP [54], and Multipath [56] applied 2D convolutions to implicitly capture interaction within the kernel size. Compared to sequence processing-based approaches, higher importance is assigned to spatial interaction.
GNNs and Graph-Attention explicitly model interactions between individual agents by combining multiple agents’ features [138] and processing them with graph convolution operators [139, 73, 140, 141] or graph attention [142, 143, 144, 145] to aggregate information. Prediction networks like VectorNet [74], LaneGCN [41], TNT [75], HOME [104], DenseTNT [106], GOHOME [84], Multipath++ [38], LaFormer [111], and BAT [114] are examples of this development.
Transformers [83] have been widely adopted in prediction: InteractionTransformer [146], mmTransformer [76], AgentFormer [105], SceneTransformer [6], HiVT [108], Wayformer [109], Motion Transformer [110], MotionDiffuser [112], Trajeglish [113] and many further works [147, 148, 149, 150, 110, 151, 152, 153, 154, 155, 156] exploit the global receptive field and attention mechanism of Transformers. All SVs can be predicted simultaneously [76], and the impact of vehicles on each others’ behaviors across different timesteps can be modeled with attention layers [105]. Attention layers can be applied along the spatial and temporal dimension jointly [105], or individually [157]. The latter can be divided into sequential [158, 159] and interleaved [6, 109] architectures. Experiments with Wayformer [109] and SceneTransformer [6] show that joint and interleaved processing outperform sequential approaches. As, joint processing causes higher computational burden, the interleaved paradigm balances best between performance and computational demand.
The timeline in Fig. 3 reflects how interaction models changed together with the scene representation. After CNNs and RNNs were applied to rasterized scene representations in early works, GNNs, attention, and Transformers took over since the advent of sparse representations.
Across all architectures, the prevalent concept is to extend interaction from a local to a global scale. HiVT’s ablation study [108] shows that both, local and global attention, are beneficial for prediction and that hierarchical attention yields the best performance.
3.3 Trajectory Decoding
The final step of prediction is to generate trajectories from latent features. As introduced in Sec. 2, this step models a probability distribution $P_{\textrm{pred}}=P(\overline{\textbf{Y}}|\overline{\textbf{X}},I)$ over possible future trajectories $\overline{\textbf{Y}}$ given observation $\overline{\textbf{X}}$ and scene context $I$. Here, we focus on two major aspects of trajectory decoding: The modeling of multimodality and the decoding principle.
Multimodality. The intention of SVs is inherently uncertain, making $P_{\textrm{pred}}$ multimodal [160]. Therefore, $P_{\textrm{pred}}$ is commonly either expressed as an explicit density function or estimated with a discrete trajectory set.
Explicit density functions. An intuitive approach to prediction is to explicitly model the probability density function of $P_{\textrm{pred}}$ [61, 66, 69, 70]. Continuous distributions like bi-variate Gaussian distributions [161], Gaussian Mixture Models [110, 162, 132], or discrete rasterized heatmaps [104, 84] can be found in many prediction works. Alternatively, object-agnostic representations like occupancy maps [53, 163], or flow fields [164, 165, 70] can be used. These representations handle multimodality implicitly [61] and were shown to be more robust against perturbations [62]. However, since no discrete trajectories are decoded, it is difficult to compare such representations with expert logs for performance evaluation.
Discrete trajectory sets are either created by sampling from an intermediate distribution or by the model design. To sample a diverse trajectory set from an intermediate distribution, generative models such as Generative Adversarial Networks (GANs) [166], Conditional Variational Autoencoders (CVAEs) [167], normalizing flows [168] or denoising diffusion [112], can be used [8, 89, 169, 170, 171]. Alternatively, recent methods apply several prediction heads to the features extracted by a common backbone to decode a set of diverse trajectories [38, 41, 172, 109]. Other ways to obtain diverse predictions are training loss functions [75], entropy maximization [39, 173], variance-based non-maximum suppression [174], greedy goal sampling [55, 97], divide and conquer strategy [175], evenly spaced goal states [176], or with pre-defined anchor trajectories [56, 177]. Nevertheless, a guarantee of coverage can only be given by specifying high-level behaviors [124] or regions of the map that must be covered by at least one prediction [76].
Decoding principles. We classify the decoding principles into one-shot, two-shot, and n-shot methods as depicted in Fig. 4. While it is a common practice in planning to decode trajectories as a sequence of actions that translate into Cartesian space using a kinematic model (cf. Sec. 4.2), only a few works do so in prediction [178, 179]. Instead, future waypoints are predicted directly from latent features [41, 38, 111, 109, 180].
Single-shot approaches are split into trajectory regression and anchor trajectory classification.
Trajectory regression directly decodes the latent features to a trajectory using a NN, typically an MLP [54, 74, 41, 38, 6, 107, 108, 109]. Especially Transformer-based architectures use direct trajectory regression decoding (cf. Fig. 3). However, these methods are prone to predict trajectories that leave the road, cut corners, or are kinematically infeasible [97]. Specifically designed loss functions for offroad driving [181, 182, 39, 183, 184], driving direction compliance [185] or distance to the centerline [186] can alleviate the problem.
Anchor trajectory classification states prediction as a classification task. Thus, a probability is estimated for a set of pre-defined trajectories [55, 177]. Defining anchor trajectories can ensure the feasibility of the prediction and allows to impose hard kinematic constraints. However, rare trajectories might not be covered by the anchors.
Two-shot methods structure trajectory decoding into meaningful subtasks that increase explainability and are more flexible than anchor trajectory classification. They comprise endpoint and completion as well as initial trajectory and refine strategies.
Endpoint and completion methods are based on the assumption that a trajectory is mostly defined by its endpoint [25]. The endpoint can either be regressed [104][84] or classified from a pre-defined set [75][106][97]. Usually, classification-based methods additionally regress an offset of the classified endpoint [75]. HOME [104] estimates the endpoint probability distribution as a rasterized map, samples potential endpoints, and then regresses the trajectory using an MLP.
Initial trajectory and refine approaches combine the ideas of anchor trajectory classification and endpoint and completion. Instead of an endpoint, an initial trajectory is classified or generated. Then, a refinement step adapts the trajectory, for example, by offset regression for each predicted waypoint [102, 56, 111]. TPNet [187] and DCMS [188] first regress an endpoint, then generate trajectory proposals and finally refine them.
N-shot methods are either initial trajectory and refine or autoregressive formulations.
Initial trajectory and refine strategies in their n-shot variants are similar to the two-shot version. The only difference lies in the refinement step, which is realized as a recurrent optimization [8, 112]. For example, MotionDiffuser [112] samples a noisy trajectory and repeatedly applies denoising diffusion to refine the trajectory.
Autoregressive approaches use recurrent decoding to perform a stepwise roll-out into the future. Iteratively, the next waypoint is predicted, and information on the resulting scene is added to the latent features [103, 189, 40, 76, 105].
Autoregressive decoding can help to enforce high-quality social interactions and scene understanding [97] but might lead to compounding errors since predicted information is used to make further predictions [190]. Further, the computational effort and computation time of these methods are usually higher than for one-shot or two-shot methods.
Fig. 3 shows the development of trajectory decoding using the categories defined in this Section. While endpoint and complete decoding gained popularity, auto-regressive formulations were recently used less often, assumingly due to computation time deficiencies. Trajectory regression decoding was always used but became prevalent with the rise of Transformer-based interaction modeling in prediction.
Figure 4: Principles of trajectory decoding.
3.4 Benchmarks
Various prediction datasets provide training, validation, and test splits for benchmarking. The most prominent ones are nuScenes [191], WOMD [192], Argoverse [193], Argoverse2 [194], inD [195], highD [196], rounD [197], exiD [198], and Interaction [199]. Evaluation is carried out by comparing the predicted trajectory of an SV to its ground truth future trajectory. To account for the inherent multimodality benchmarks are commonly based on a winner-takes-all evaluation. This means the model predicts a fixed number of output trajectories, e.g., six [194] or ten [191], and the evaluation takes only the best one among them into account. Alternatively, predictors that output a parametric distribution are often evaluated based on the Negative log likelihood (NLL) of the ground truth under the predicted distribution. This assessment also factors in the uncertainty predicted by the model and evaluates the whole distribution instead of only a single trajectory.
4 Planning
The planning task is to find a trajectory $Y_{\textrm{EV}}$ for the EV that factors in safety, comfort, kinematic feasibility, and goal-directedness based on the observations $X_{\textrm{EV}}$ and $X_{\textrm{SV}}$ as well as additional context $I$ and optionally $\overline{\textbf{Y}}_{\textrm{SV}}$ (cf. Sec. 2). In this section, we will first give a comprehensive overview of common input (Sec. 4.1) and output representations (Sec. 4.2). Then, we address goal-conditioning in Sec. 4.3 before we categorize existing works and discuss common paradigms in Sec. 4.4. Finally, we briefly describe existing benchmarks in Sec. 4.5.
4.1 Input Representation
We classify planning input representations into interpretable intermediate representations and latent features.
Interpretable intermediate representations such as rasterized or sparse scene representations (that were discussed thoroughly in Sec. 3.1) are common inputs to planning components in modular ADSs. Thus, many works exist that build on rasterized [58, 67, 124, 200] or sparse inputs [125, 131, 201, 130, 24]. Similar to prediction, a clear trend can be observed from rasters to sparse inputs.
Latent features. E2E ADSs directly map sensor information to actions (cf. Fig. 1[9, 202, 20]. Thus, instead of $X_{\textrm{EV}}$, $\overline{\textbf{X}}_{\textrm{SV}}$, and $I$, the planner’s inputs are latent features.
The benefit of using learned latent intermediate representations is that no manual engineering effort is needed to design the interfaces between different modules. The main drawback of hand-engineered interfaces is that they have to apply to the long-tailed distribution of potential driving scenarios, and every piece of information discarded by the interface cannot be recovered. For instance, if a vehicle is currently occluded, but its presence can be inferred from the reflection of its headlight in another car, bounding box representations [58], occupancy fields [70], or affordances [115, 119] will not be able to propagate this information to downstream modules [203, 133]. In contrast, latent feature representations allow uncertainty propagation, enabling downstream modules to compensate for errors at earlier stages, such as undetected vehicles [61].
The main disadvantage of latent feature representations is the lack of interpretability. In case of failures, evaluating which part of the system caused the error is extremely difficult. This aggravates tuning and debugging of pure E2E systems. Interpretable E2E systems seek to compensate for this by producing additional intermediate representations [66]. In contrast to modular stacks, these interpretable intermediate outputs are not used by the planner. Instead, they are only used for additional supervision and model introspection.
The timeline in Fig. 3 reveals that interpretable input representations are becoming increasingly popular compared to latent input features.
4.2 Output representation
The planning output is typically represented as a sequence of future states or control actions.
Future states are given by a sequence of future 2D position and heading angles representing the planned trajectory of the EV over time [22, 62, 58]. This trajectory is then passed to a downstream controller tasked with tracking it [22]. The trajectory representation is well interpretable and can be easily checked for collisions, traffic law violations, or deviations from the driveable surface [129, 68]. Moreover, it is agnostic of the kinematic parameters of the controlled vehicle, enabling generalization to other vehicles. However, the ultimate driving performance also depends on the downstream controller and how well the planner and controller are tuned for each other. Validating and testing the system can require iterations between both components, which slows down development.
Future actions are an alternative representation of planned behavior. This is mostly adopted by E2E ADSs that regress acceleration and steering angle without the intermediate trajectory representation [117, 118, 119]. Besides ensuring kinematic feasibility, it can promote comfort as it is directly related to the magnitude of the actions.
However, the resulting behavior depends on the controlled vehicle’s individual dynamics [204, 19]. Hence, generalization to other vehicles is limited.
Fig. 3 depicts that early E2E ADSs planned in actions, while trajectory outputs recently gained popularity.
4.3 Goal Conditioning
The overall task of ADSs is to navigate safely toward a destination. Hence, goal-directedness is one main criterion that determines the suitability of a potential plan. Lane-level route information can be provided to the planner by a navigation system. Popular benchmarks provide route information through a high-level command such as ’turn left’ or ’go straight’ [205], through sparse goal points along the route [205] or by annotating lanes within the map as on-route or off-route [206]. For example, in the widely used Carla simulator [205], goal locations are sparsely sampled along the route, and in each planning step, the closest goal location is provided to the planner. We define four categories to describe how this goal information is incorporated into the planning algorithm: input features, separate submodules, routing cost, and route attention. While the first two were originally introduced in [117], the latter were proposed more recently [201, 207, 126, 208].
Input features are the most straightforward way to incorporate goal information and have been widely adopted [120, 20, 124, 209, 210, 24, 63, 64, 118, 23, 122]. Different methods exist to pass this input to the planning module depending on how the route information is represented. If the lanes within the map information are annotated as on-route or off-route, then this information can be used as a separate semantic channel in raster maps [58]. Similarly, vector inputs describing the lane centerlines can have an additional flag reflecting this information [201].
High-level commands were used as input feature in LBC [124], LAV [67], and AD-MLP [208]. It was found that inputting high-level intention repeatedly at different stages of the network increases its generalization capability [23]. Transfuser [20] leverages sparse goal locations as 2D features in the final trajectory decoding step.
While using goal information as an input feature is simple, goal compliance is not guaranteed. On the contrary, [201, 117] demonstrated that the model could ignore this additional input and instead rely on other potentially spurious correlations, such as between the current and the future kinematic state.
Separate submodules are used only with high-level commands, which serve as a switch between command-specific submodules [117, 70]. Balancing the dataset to prevent it from being dominated by a single driving mode, such as lane following, is avoided in this setup, as each block is only trained on a single high-level command. However, this approach requires a fixed number of high-level commands to be defined in advance.
Routing cost. A third option is to plan trajectories by optimizing a cost term for routing [66, 69, 62]. Such cost functions can measure the remaining distance to a goal location [62], the progress along or distance to an on-route lane [66]. Goal-directedness is balanced with further competing objectives such as safety and comfort. Thus, the planner can flexibly react to difficult situations, e.g., it could drive off the road to prevent a collision. However, tradeoffs between safety and comfort versus progress can be undesirable.
Route attention applies spatial attention to relevant parts of the map, i.e., the route. This can be achieved by removing the off-route portion of the map from the input features, e.g., PDM-Open [126] and PEP [94] discard almost the entire map information and only keep points along a lane centerline on the route. Similarly, GC-PGP [201] removes off-route nodes from the lane graph after information aggregation and TPP [134] and DTPP [136] discard all EV trajectory proposals that do not comply with the route.
Fig. 3 shows that early models did not condition their plan on a goal at all. Afterward, submodules and input features dominated. While using input features is easy to implement, it is less effective, whereas submodules represent a major architectural constraint. The most interpretable method is using a routing cost function. Recent models introduced route attention, which is more flexible. Both route attention and routing cost allow to condition on goals that are not reflected in the training dataset, thus improving generalization.
4.4 Planning Paradigms
Recall that in Eqn. 5 we defined planning to be a function $f$ mapping from observational inputs $X_{\textrm{EV}}$, $\overline{\textbf{X}}_{\textrm{SV}}$, $I$ and predictions $\overline{\textbf{Y}}_{\textrm{SV}}$ to a trajectory $Y_{\textrm{EV}}$. In the following, we focus on the planning function $f$. To this end, we decompose $f$ into two parts: A proposal generator $g$, yielding a set of multiple potentially suitable trajectories $\hat{Y}^{(i)}_{\textrm{EV}},i=1,..,N_{\textrm{proposals}}$ and a proposal selector $h$ that selects the final plan $Y_{\textrm{EV}}$ among the proposed set. Hence, in this work, we propose to represent the planning function $f$ as
$f=h(g(X_{\textrm{EV}},\overline{\textbf{X}}_{\textrm{SV}},I,\overline{\textbf{Y}}_{\textrm{SV}})\,.$ (7)
Based on this, we distinguish three different paradigms adopted for the planning task: Cost function optimization, regression, and hybrid planning.
Cost function optimization. As planning aims to find a trajectory that optimizes objectives such as safety, comfort, and progress, it is a natural approach to design and optimize a cost function that balances these potentially conflicting objectives. Cost function-based planning methods [121, 61, 211] rely entirely on the selection function $h$ to find a suitable trajectory, i.e.,
$h=\operatorname*{argmin}_{i}c\left(\hat{Y}^{(i)}_{\textrm{EV}}\right),$ (8)
where $c$ is a cost function of the planned trajectory. Consequently, the proposal generator $g$ may randomly sample feasible motion profiles [61] or obtain them by clustering real-world expert demonstrations [70]. Traditional non-learning-based methods rely on hand-crafted cost functions [212, 213, 214, 215, 216, 217, 218, 123]. However, designing a cost function that effectively balances these goals and generalizes to the long-tailed distribution of possible driving scenarios is challenging. Learning-based methods aim to address this by learning a cost function directly from expert demonstrations [219, 220, 103]. Similarly, [119, 115] learn to regress driving-relevant affordances such as the distance to the leading vehicle or the heading angle relative to the lane, based on which a downstream controller infers driving commands.
The learned cost function can be non-parametric [61, 221, 222] or comprise hand-designed sub-costs of which only the weights are learned [69, 70, 66, 121]. The non-parametric cost function of the NMP [61] assigns a cost to each potential future EV position within the planning horizon and is represented as a 3D-BEV-Tensor. In contrast, hand-designed sub-costs commonly evaluate safety, comfort, traffic rule compliance, and progress [121, 70, 66].
Regression approaches rely entirely on the proposal generator $g$. The selection function $h$ is the identity, and all burden is placed on the proposal generator $g$, which only generates a single proposal $Y_{\textrm{EV}}=\hat{Y}^{(1)}_{\textrm{EV}}$. E2E methods fall into this category. Regression approaches can learn to directly predict future waypoints [201, 20, 24] or actions [9] by supervised learning [223]. This form of imitation learning, i.e., supervised learning from expert data, is called behavior cloning (BC) [224, 190, 225]. This approach makes the manual design of the cost function obsolete. However, BC models are vulnerable to distributional shifts, which occur when the model reaches states not covered by the training data. Besides, regressed trajectories are generally not guaranteed to be kinematically feasible. UrbanDriver [24] counters this by training the model directly in a closed-loop simulation. However, this requires the entire simulator to be differentiable. CW-ERM [226] rolls out the policy in a non-differentiable simulation to identify important training scenarios and assign sample weights.
Hybrid planning describes methods that combine both ideas. First, a set of candidate trajectories is regressed by the proposal generator $g$, then the best one is selected with a cost function in $h$ [227]. For instance, [130, 131, 228, 229, 132] perform collision checks on the proposed trajectories to rule out unsafe proposals. In this setting, it often remains unclear which parts of the overall objective are enforced by the proposal generator and which by the cost function.
Fig. 3 shows that regression is the dominating planning paradigm for E2E systems. Cost function-based planning is mainly used in the work of researchers from Uber ATG and the University of Toronto [68, 66, 69, 70].
4.5 Benchmarks
Different methods exist to evaluate planning approaches, namely open-loop evaluation and closed-loop simulation.
Open-loop evaluation is similar to an evaluation in the prediction task. It compares the planner output $Y_{\textrm{EV}}$ to that of an expert planner $Y_{\textrm{GT}}$ [230]. However, this does not give the planner control over the EV and ignores the problems arising from compounding errors and distributional shifts [231]. More severely, recent work demonstrated that open-loop evaluation results do not correlate with driving performance [126, 232].
Closed-loop simulation in contrast, lets the planner control the EV. Instead of merely comparing the planner output $Y_{\textrm{EV}}$ to that of an expert planner $Y_{\textrm{GT}}$, the evaluation is based on metrics for comfort, safety, and progress. Thus, closed-loop evaluation relates better to real-world driving but requires a simulator to let the model interact with its environment. Prominent simulators are Carla [205], nuPlan [206], Waymax [233], Highway-Env [234], and CommonRoad [235]. Carla is a high-fidelity simulator that involves sensor simulation, whereas nuPlan, Waymax, CommonRoad, and Highway-Env operate on abstract object representations (e.g., bounding boxes) without rendering camera or lidar sensors. Many benchmark datasets were proposed for the Carla simulator [232, 236, 237, 20]. However, the scenarios are purely synthetic, which questions the generalizability to the real world. In contrast, nuPlan and Waymax follow a data-driven approach, which uses recordings from real-world scenarios to initialize the simulation. Thus, users have access to corresponding large-scale real-world datasets to train their method before deploying it in simulation. CommonRoad is compatible with several smaller datasets [195, 196, 197, 198, 199]. Highway-env is based on simplistic synthetic environments.
When simulating regular driving scenes, rare and critical scenarios are naturally under-represented. This problem can be tackled with data curation [238, 239] or additional test case generation. These can be generated manually [240, 241, 242, 243, 244, 245] or automatically from recorded driving logs [246, 247, 248]. Other methods generate adversarial critical scenarios by augmenting actors within the scene [249, 250, 251, 252, 253, 254, 255, 256] or the scene itself [257, 258, 259, 260, 261]. Similarly, [5, 64] investigate human-robot interaction more closely with hand-crafted scenarios – an aspect that is uncared for in average driving situations.
Another major challenge in closed-loop evaluation is to realistically simulate the SVs. Simple simulators such as Highway-Env [262] rely on non-reactive agents that follow predefined trajectories [263]. In the non-reactive version of nuPlan, and Waymax, these trajectories are obtained from real-world recordings. However, if the EV behaves differently than the expert during recording, this can result in collisions caused by the SVs, as they do not react to the EV. Hence, this simulation is usually limited to short sequences, where the EV stays close to the expert’s trajectory [264]. Carla, nuPlan, and Waymax also provide a reactive driver model for the SV, such as the Intelligent Driver Model (IDM) [16, 17, 265, 266, 267, 268]. Another line of work focuses on learning-based methods to simulate realistic driving behavior [269, 270, 271, 272, 273, 274]. In Waymax [233], users can plug in their own model to control objects in the scene.
5 Integrating prediction and planning
Figure 5: Connection between the integration principles, ego-to-agent relationship, and cost function. Red arrows for the robot leader planning indicate that applying these cost function categories would be possible but lacks a thorough empirical evaluation. We highlight this as an interesting direction for future research.
After reviewing key aspects such as interfaces and architecture of planning and prediction components found in ADSs, we change perspective and discuss the properties of the IPPS. In particular, we will highlight the influence of design decisions in the IPPS on behavior in interactive scenarios. We will start by categorizing existing works based on which components the IPPS comprises (Sec. 5.1). Then, we will describe the implications for the interactive behavior which follow from design choices in modular integrated systems (Sec. 5.2). Subsequently, different concepts of safety and contingency in modular integrated systems are discussed in Sec. 5.3. Finally, Sec. 5.4 discusses possible combinations of these categories. An overview is provided in Fig. 5.
5.1 Integration Principles
In this work, we distinguish among three design principles for IPPS. An overview can be seen in Fig. 6.
Monolithic E2E planning describes methods where the IPPS consists of a single E2E planner, which maps the state inputs $\overline{\textbf{X}}_{\textrm{SV}}$, $X_{\textrm{EV}}$, and $I$ to the EV’s trajectory $Y_{\textrm{EV}}$. Thus, the SVs’ future behavior and the interactions among them, as well as between the SVs and the EV, are not modeled explicitly. However, the expert’s actions, which the E2E planner learns to imitate in training, are based on reasoning of how the future will unfold. By learning to imitate actions, the model could exhibit behavior that is implicitly based on such reasoning as well.
For instance, consider a scenario where the EV’s leading vehicle brakes. An expert or a human driver will base their driving decision on the likely future scenario that the safety margin will shrink and thus brake as well. Hence, an E2E planner that imitates the expert’s behavior makes a decision that is implicitly based on a predicted future.
Pioneered by Pomerlau [9], there is a vast body of literature on IPPS based on monolithic E2E planners [19, 116, 117, 118, 22, 62, 10, 63, 64, 23, 67, 122, 236, 24]. The main disadvantage is their black-box nature, which makes model introspection and safety verification difficult.
Interpretable E2E planning frames prediction as an auxiliary learning task [58, 61, 120, 125]. Explicit predictions are an additional model output trained jointly with the planning task. Usually, a shared backbone encodes features that are used for both tasks. Then, separate heads decode the respective output representations.
The additional objective provides a learning signal to the backbone. It acts as a regularization which can make the learning process more sample-efficient and can increase generalization capabilities [58, 124, 119] as it was previously demonstrated for other auxiliary tasks in prediction and planning like speed prediction [64], semantic mapping [275], and occupancy forecasting [66]. However, the additional learning objective results in a trade-off between the different tasks, usually resulting in a hyperparameter that balances the respective losses and has to be tuned empirically.
Compared to monolithic E2E planning, the explicit prediction adds interpretability and facilitates introspection. Nonetheless, both IPPS designs cannot provide safety guarantees. Hence, we attribute the increased driving performance to the regularizing effect of the additional objective.
Manual integration of a planner with a prediction module means that a separate subsystem is employed for each of the tasks. The interplay of both tasks is manually engineered based on domain knowledge. The most widely adopted integration is a sequential pipeline, where first the prediction module is inferred, and its output is fed to a subsequent planning module [70, 66, 69, 123]. Since this design is unable to reflect the impact of the EV’s plan on the SVs, an alternative is to swap the order of prediction and planning: First, candidate plans are generated, followed by a conditional prediction module. Then, the best plan for the EV is selected among the candidates [129, 134]. Alternative methods iterate between prediction and planning [276] or do both jointly [132]. In the following, we give a chronological overview of different models and highlight the integration design of prediction and planning.
(a) Monolithic E2E IPSS
(b) Interpretable E2E IPSS
(c) Manually Integrated IPSS
Figure 6: IPPS design principles that show how prediction and planning can be integrated. Monolithic E2E methods (a) use no explicit prediction, while interpretable E2E methods (b) use it as a regularization in training. Manually integrated IPPSs (c) combine a modular prediction with a modular planning method using hand-designed interfaces.
PRECOG [65] uses a probabilistic model for conditional multi-agent forecasting. It implements autoregressive trajectory decoding for EV and SVs jointly. At each time step, the state of all agents is updated and used as input for all other agents in the next update step.
PiP [129, 277] reverses the view of IPPS. Its core idea is that SVs behave differently depending on the EV’s actions. A trajectory generator provides a set of candidate plans independent of the SVs. Each candidate plan is then used to condition prediction. Finally, the optimal plan is selected using a cost function.
DSDNet [68] combines E2E and modular model design. Each of the sequential NN modules additionally has access to the high-dimensional features of the perception backbone, complying with the hypothesis [20] that scene context is important for robust planning. For each detected vehicle, a set of potential future trajectories is predicted. To select the plan for the EV from a sampled candidate set, a hand-crafted cost function is applied, which quantifies the collision probability of the EV candidate trajectories and the SVs’ predicted trajectories.
P3 [66], LookOut [69], and MP3 [70] use cost functions to link prediction and planning. P3 predicts an occupancy map, LookOut [69] employs predicted trajectories to evaluate different pre-defined EV candidate trajectories. MP3 [70] extends P3 into a mapless approach by predicting an online map. A ’dynamic state map’ comprising detections and prediction is used together with the intended route to evaluate potential trajectories.
SafetyNet [130] parallelly applies a monolithic E2E planner with implicit prediction and an explicit prediction module. Multiple collision checks between EV plan and SVs predictions are performed. If the initial plan is classified as unsafe, a hand-crafted fallback layer generates a lane-aligned trajectory. SafePathNet[131] improves upon SafetyNet [130] with a transformer module for joint prediction and planning. In a mixture-of-experts approach, multiple modes are predicted and ranked. The ranking is updated by checking for collisions with the most probable SVs’ predictions. Finally, the highest scoring EV mode is selected.
DIPP [7] jointly predicts the trajectories for all agents, including the EV, which is treated as an SV. The mode with the highest probability is selected for each vehicle. A differentiable nonlinear motion planner then performs iterative, local optimization under consideration of a kinematic model and a hand-crafted cost function.
Inspired by hierarchical game-theoretic frameworks [278, 279, 280, 281], GameFormer [132] models interaction between agents as a level-$k$ game [282, 283]. Thus, a transformer decoder models the interaction between all agents, including the EV, by iteratively updating the individual predictions based on the predicted behavior of all agents from the previous level. Hence, the model yields a joint prediction for all agents, including the EV.
UniAD [133] implements a modular system that is trained E2E based on the planning objective. Queries are used as the interface between the submodules so that the planning module can attend to agent-level features from previous network layers. Various works adapted or extended this approach: VAD [135] and Hydra-MDP [284] use the output of previous layers explicitly for collision and map compliance checks. PPAD [276] repeatedly iterates between prediction and planning based on the perception modules’ outputs and GenAD [285] adapts a generative decoder to regress SV and EV trajectories jointly.
TPP [134], DTPP [136], and MBAPPE [286] represent multiple future outcomes in scenario trees, where branches reflect combinations of SV predictions and EV plans. The best plan is obtained through a tree search.
Compared to other integration principles, manual integration requires more engineering effort but restricts the solution space in a meaningful way by incorporating prior knowledge. Often, manually integrated IPPS provide a high explainability and safer plans than E2E systems [287, 131].
The timelines in Fig. 3 show how early works relied mostly on implicit predictions in monolithic E2E designs. Starting with ChauffeurNet [58] and PRECOG [65], explicit predictions became more popular. More recently, manual integration of E2E differentiable modular components gained popularity [203, 133].
5.2 Ego-to-Agents Relationship
After reviewing the integration principles from a system architecture perspective, in the following, we will analyze the manual integration category from a different perspective, namely the intended ego-to-agents relationship. This is of special importance in highly interactive scenarios since the EV needs to base its driving decisions on the observed and expected behavior of surrounding agents. However, as [5] pointed out, the EV also needs to be aware that it can influence the behavior of others.
For instance, in the example in Fig. 7, the plan of the EV can also influence the behavior of the SV: If it approaches the intersection faster, the SV might decelerate and yield to the EV. To describe if and in which way the mutual interaction of the EV and surrounding traffic is taken into account in the planning step, [5] introduced the following four categories: robot leader planning, human leader planning, co-leader planning, and joint planning. In this context, human refers to the surrounding traffic, and robot refers to the self-driving vehicle. In the following, we briefly describe these categories and discuss existing work.
Robot leader planning. The EV plan is inferred based on the current state, and the prediction step for the whole environment is conditioned on it. Hence, the EV can effectively seek to make the SV react to its trajectory [88, 128, 127]. This can result in aggressive driving behavior. For instance, in the example in Fig. 7, the EV would reason that when it follows the fast-progressing plan, the observed SV will yield to it to prevent a collision.
Human leader planning is the opposite of robot leader planning. The plan is based on the forecasted behavior of SVs [61, 62, 70, 66]. Hence, the EV shows a reactive behavior. It does not model the influence of the EV’s plan on the SVs, which can lead to underconfident behavior [5]. This is the case for IPPS that predict and plan sequentially. In the unprotected right turn example (cf. Fig. 7), the EV tries to find a plan that is suitable given the two predicted SV behaviors without being aware that it can influence them. Consequently, it will favor the slower plan.
Joint planning describes systems that are aware that vehicles (including ego) interact with each other. The EV’s plan is obtained by global optimization over all agents [288]. Hence, the IPPS system deterministically approximates a joint objective [289] under the assumption that an optimal outcome exists. Hence, it erroneously assumes that the behavior model used for SVs is exactly known and that every traffic participant optimizes this same global objective. For example, if the EV in Fig. 7 squeezes in before the approaching SV, this may be optimal with respect to a reasonable global objective. Nonetheless, there is no guarantee that the SV will behave accordingly. Consequently, [5] demonstrates how this can lead to fatal errors.
Co-leader planning is the most complex paradigm. It models the influence of SVs’ potential future behaviors on the EV plan as well as their reaction to potential ego trajectories [290, 291, 5]. In contrast to the joint planning category, the surrounding agents’ behavior is not assumed to be deterministic. The EV has to consider this uncertainty in the planning step, e.g., by being prepared for multiple future outcomes and reacting accordingly. Additionally, co-leader planning expects the SV to make plans for different potential behaviors of the EV. The EV can exploit this by behaving in a way that makes the SVs reveal their intention, which in turn lets the EV make more confident plans. For instance, in the example in Fig. 7, the EV can try to slowly approach the intersection to provoke a reaction of the SV. This reaction can either indicate that the SV is going to make a turn, e.g., with a turn signal or by slowing down, or the opposite, i.e., accelerating, to indicate that the SV will continue straight. The former will enable the EV to confidently decide in favor of the faster-progressing plan. CfO [5] originally introduced this paradigm, leveraging normalizing flows to represent multi-agent behavior as a learned coupling of independent variables. However, the intricate nature of this architecture may pose obstacles to wider adoption. Furthermore, their benchmark is simplistic, encompassing only a few scenarios with hard-coded agent behaviors. TPP [134] presents a readily comprehensible tree-structured framework that implements co-leader planning. It involves sampling kinematically feasible trajectories for the EV and then conditioning prediction on each of them. An autoregressive formulation for prediction ensures causality: It can be seen as a way to roll out the scene with SV predictions and the EV plan such that both can only access the observation history up until the respective rollout step. Finally, the best EV trajectory proposal is selected with a dynamic programming algorithm. DTPP [136] extends this with a transformer-based conditional prediction module, which allows the forecast of all agents conditioned on multiple EV plans at once. It also replaced the dynamic program with a learnable cost function and applied node pruning to reduce computational burden. Still, the EV reaction to the prediction is taken over entirely by the cost function because the sampled ego-proposals are non-reactive with respect to the SV predictions.
Fig. 3 shows that no clear trend toward one of the paradigms can be observed. While Co-leader planning is theoretically the most capable paradigm, it has not yet been shown empirically if one of these four theoretical concepts is superior to the others and how it can be realized in the design of system architectures. This may also be related to a lack of comprehensive empirical benchmarking focused on interactive behavior.
5.3 Safety and Contingency
Figure 7: Demonstration of contingency on an unprotected right turn. Let’s assume two expected future behaviors of the observed SV, namely $Y^{(1)}_{\textrm{SV}}$ and $Y^{(2)}_{\textrm{SV}}$ with probabilities of $P(Y^{(1)}_{\textrm{SV}})=0.2$ and $P(Y^{(2)}_{\textrm{SV}})=0.8$. Two candidate plans for the EV are considered $Y^{(1)}_{\textrm{EV}}$ (purple) and $Y^{(2)}_{\textrm{EV}}$ (orange). While the purple plan is only safe if the unlikely prediction $Y^{(1)}_{\textrm{SV}}$ occurs, the orange plan is safe under all predictions. Thus, it is preferred under worst-case assumptions. When planning with marginalized predictions, the cost function has to balance between the high progress of the purple plan and the potential safety risk if the unlikely prediction $Y^{(1)}_{\textrm{SV}}$ occurs. A contingency planner could exploit that both candidate plans have a common short-term action. Hence, it could opt for the purple plan as it is aware that it could still re-plan and fall back to the orange one.
The ego-to-agents relationship categories discussed above highlighted that considering multiple potential future scenarios is essential to safety and contingency planning. In the following, we will discuss how this can be incorporated into the cost function component $h$ that we defined as part of the planning function $f=h(g(X_{\textrm{EV}},\overline{\textbf{X}}_{\textrm{SV}},I)$ (cf. Sec. 4.4). We will analyze the implications of different design choices on safety and contingency. To this end, we introduce the following notation: We denote the number of relevant future scenarios with $N_{s}$. These scenarios can be reflected by a multinomial distribution $P\left(\overline{\textbf{Y}}^{(i)}_{\textrm{SV}}\right),i=1,..,N_{s}$ describing the behavior of SVs in each scenario. Based on how this distribution is reflected in the design of the IPPS, we form the following three groups of existing methods regarding safety and contingency: Planning with marginalized predictions, worst-case planning, and contingency planning. They are briefly outlined in Fig. 7.
Planning with marginalized predictions describes IPPS that do not explicitly distinguish multiple future scenarios $N_{s}$. This means that the prediction $\overline{\textbf{Y}}_{\textrm{SV}}$ is marginalized over the future outcomes $\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}$. This can be done explicitly, i.e., by integrating over the scenario probabilities $P\left(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}\right)$
$\overline{\textbf{Y}}_{\textrm{SV}}=\sum_{j}^{N_{s}}\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}\cdot P\left(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}\right)$ (9)
or implicitly, i.e., by representing all possible scenarios in a joint representation instead of differentiating scenarios. For instance, IPPS that incorporate the prediction in their cost function $h$ and do not consider separate future outcomes when evaluating a proposal fall into this category [61, 70, 66]. Hence, their selection function takes the form
$h=\operatorname*{argmin}_{i}c(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}_{\textrm{SV}}).$ (10)
Similarly, this is the case for all systems in the monolithic E2E category (cf. Sec. 5.1[201, 20, 24, 9, 19], as they do not explicitly consider multiple future scenarios and hence – if at all – learn marginalized predictions. Broadly speaking, this means that the IPPS assumes that all $N_{s}$ scenarios will occur to a certain degree at the same time. As can be seen in the example of Fig. 7, the cost function needs to trade off unlikely but dangerous scenarios (e.g. collision) against very likely low-cost scenarios. This is safety-critical since planners need to be cautious, especially w.r.t. unlikely but dangerous events.
Worst-case planning refers to IPPS that are aware that multiple future outcomes $\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}$ exist. In this category, no probabilities $P(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})$ are considered. Instead, each proposal is evaluated based on the worst-case scenario. The motivation behind this assumption is that the final output trajectory must be safe in each possible scenario. Consequently, the selection function $h$ can be formulated as
$h=\operatorname*{argmin}_{i}\max_{j}c(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}).$ (11)
This paradigm strongly focuses on collision avoidance and can result in overly conservative behavior, as shown in Fig. 7. While this assumption is not a popular choice in DL-based prediction and planning, it is broadly applied for rule-based safety layers such as RSS [292] and can be seen as an intermediate step between planning with marginalized predictions and contingency planning.
Contingency planning prepares for the unknown future development of the scene by taking into account different future scenarios $\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}$ and their probabilities $P(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})$. The resulting plan hedges against worst-case risks while enabling progress in expectation. Both the cost function-based and the hybrid planning paradigm can have these properties, and the selection function $h$ takes the general form
$h=f_{c}(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}^{(j)}_{\textrm{SV}},P(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}))\,.$ (12)
where $f_{c}$ is a contingent selection function.
For instance, [69, 287, 293] find a plan that is safe for scenarios on the short horizon (independently of their probability, i.e., worst-case) and also optimize the expected cost on the long horizon (marginalization). Thus,
$h=\operatorname*{argmin}_{i}\left(\max_{j}c_{s}(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})+\mathbf{E}_{j}[c_{l}(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})]\right),$ (13)
where $c_{s}$ and $c_{l}$ consider the short-term and the long-term part of the proposal $Y^{(i)}_{\textrm{EV}}$, respectively. In the example in Fig. 7, this planner allows the most progress while still ensuring safety. CfO [5] further differentiates between passive contingency where the EV is contingent upon the SVs [69, 287, 293, 294] and active contingency where the EV also expects the SVs to be contingent upon it [290, 291, 295].
5.4 Possible Combinations
In the previous sections, we have described three ways to categorize the integration of prediction and planning methods, namely (1) integration principle, (2) ego-to-agents relationships, and (3) safety and contingency. In the following, we discuss possible combinations of these three dimensions. Fig. 5 shows an overview. Our key insight is that the categories we characterized describe the differences between IPPS methods on different levels. While the integration principle focuses on a high-level system architecture, the ego-to-agent relationship is mostly based on interactive behavior. The safety and contingency consideration is founded on the cost function, i.e., a specific design choice for the proposal selection. In the following, we highlight associations between these categories.
In case of monolithic E2E [9, 20, 24, 19, 117, 64, 10, 121, 23, 67, 122, 62, 22, 118, 116, 201] and interpretable E2E [58, 61, 296], no part of the system architecture explicitly reflects an intended interactive behavior. Hence, they cannot be assigned to the ego-to-agents relationship categories, and we determine their ego-to-agents relationship as unknown.
For instance, imagine that a model is trained on demonstrations from an expert who follows the co-leader planning paradigm and employs contingency planning. Given that the model excels at imitating the expert, it could be assumed that the model behaves in the same way. However, there is no guarantee for this, and it could be equally possible that the model is unable to reason effectively about SVs, resulting in a behavior similar to the robot leader paradigm.
Similarly, as E2E models do not use hand-crafted cost functions with a pre-defined structure, the safety and contingency categories described in Sec. 5.3 do not apply to them.
In contrast, inferring the ego-to-agents relationship of modular systems is often possible. Thus, we can further divide manually integrated architectures into robot-leader, human-leader, joint planning, and co-leader planning.
There is a plethora of manual integration methods following the human leader paradigm (cf. Fig. 3). By following a traditional sequential scheme of prediction and planning, [70, 66, 69, 120, 68, 130] do not model the influence of the ego plan on SVs. This can be combined with any of the three safety and contingency categories. Works like DSDNet [68], P3 [66], MP3 [70], UniAD [133], and  [297] leverage marginalized predictions.
While applying a worst-case cost function is theoretically possible, most works that can distinguish multiple future scenarios aim to make contingency plans [69, 287, 293].
Similarly, we argue that the robot leader paradigm can be combined with all three cost functions. Consider a simple robot leader model that first identifies potential ego plans and then predicts the future behavior of SVs conditioned on each plan. The final selection of the EV’s plan can be based on a cost function belonging to each of the three categories. However, existing works use specialized cost functions that do not follow the structure we outlined in Sec. 5.3, e.g.,[129, 134, 128, 88]. We want to highlight that systematically combining robot-leader architectures with corresponding cost functions could be a promising direction for future research. Especially contingent or worst-case cost functions could mitigate the inherent problem associated with robot leader planning, namely relying on SVs to react in potentially unreasonable ways for the interests of the EV.
As described in Sec. 5.2, the active contingency resulting from the co-leader planning paradigm implies that the EV makes contingency plans while being aware that SVs are contingent upon it. Hence, it naturally employs a cost function that supports contingency plans.
Finally, the joint planning paradigm relies on optimizing a joint cost function for all traffic participants in the scene. Thus, none of the safety and contingency concepts are applicable, and we denote such a cost function as joint cost.
6 Outlook and Conclusion
Based on the overview we gave on deep learning-based prediction, planning, and their integration in automated driving systems, we identify two core challenges for future research: System design and comprehensive benchmarking. After discussing both, we draw a final conclusion.
System Design. Employing a traditional, strictly sequential system consisting of perception, prediction, planning, and control is still a popular choice. Our survey suggests that this approach is not able to meet the requirements set for driving systems. Alternative methods integrate prediction and planning in a way that allows conditioning prediction on potential ego plans and vice versa. Nonetheless, it remains unclear which exact integration architecture is most effective. Especially in the increasingly popular field of interpretable end-to-end systems, it is unclear how prediction and planning should be integrated.
Comprehensive Benchmarking. We discussed different aspects of the integration of prediction and planning. However, no comprehensive empirical benchmark reproduces and analyzes their strengths and weaknesses. Such an overview would help to better understand the effects of different ego-to-agents relationships and safety/contingency paradigms. This requires simulation in realistic and highly interactive scenarios with realistic driver models for surrounding vehicles and expressive interaction metrics. Moreover, benchmarks must cover corner case scenarios and test the generalization to distributional shifts.
In this work, we surveyed and analyzed the integration of prediction and planning methods in automated driving systems based on a comprehensive overview of the individual tasks and respective methods. We described, proposed, and analyzed categories to compare integrated prediction and planning works and highlighted implications on safety and behavior. Finally, we pointed out promising directions for future research based on the identified gaps.

### v3

@@
-Rethinking the Integration of Prediction and Planning in Deep Learning-Based Automated Driving Systems: A Review
+The Integration of Prediction and Planning in Deep Learning Automated Driving Systems: A Review
 Abstract
@@
 While this accounts for the influence of surrounding traffic on the ego vehicle, it fails to anticipate the reactions of traffic participants to the ego vehicle’s behavior.
-Recent models increasingly integrate prediction and planning in a joint or interdependent step to model bi-directional interactions.
+Recent methods increasingly integrate prediction and planning in a joint or interdependent step to model bidirectional interactions.
 To date, a comprehensive overview of different integration principles is lacking.
-We systematically review state-of-the-art deep learning-based prediction and planning, and focus on integrated prediction and planning models.
-Different facets of the integration ranging from model architecture and model design to behavioral aspects are considered and related to each other.
+We systematically review state-of-the-art deep learning-based planning systems, and focus on how they integrate prediction.
+Different facets of the integration ranging from system architecture to high-level behavioral aspects are considered and related to each other.
 Moreover, we discuss the implications, strengths, and limitations of different integration principles.
@@
 Automated Driving, Motion Prediction, Motion Planning, Deep Learning
+Note: A beginner-friendly introduction to prediction, planning, and their integration in automated driving systems can be found in an earlier version of our work at https://arxiv.org/abs/2308.05731v2.
+The following version is intended for more experienced researchers in the field.
 {textblock*}
-6.6in(0.95in,0.15in) SUBMITTED TO REVIEW AND POSSIBLE PUBLICATION. COPYRIGHT WILL BE TRANSFERRED WITHOUT NOTICE.
+7in(0.75in,0.15in) (c) IEEE 2024.
 Personal use of this material is permitted.
 Permission must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
-1 Introduction
+I Introduction
 Automated Driving (AD) remains a challenging endeavour.
@@
 Fig.
-1) address prediction and planning as separate tasks.
+2) address prediction and planning as separate tasks.
 The predicted behavior of surrounding traffic agents is leveraged to plan a suitable behavior for the ego vehicle.
@@
 In fact, prediction and planning are no sequential problems and should be tightly coupled in automated driving systems [5, 6, 7].
-1.1 Scope
-In this work, we review methods to integrate prediction and planning into an automated driving system (ADS).
+Fig.
+1 highlights the interdependence of both tasks.
+Figure 1: How should the ego vehicle behave?
+Various behaviors are possible for the ego vehicle.
+Here, we depict two in red and orange.
+Their consequences depend on the behavior of the observed surrounding vehicle, which is depicted by blue arrows (prediction A or prediction B).
+At the same time, the surrounding vehicle might react to the ego vehicle’s action, i.e.
+the surrounding vehicle’s behavior also depends on the ego vehicle’s decision.
+Different methods exist to forecast the behavior of the surrounding vehicle, and various ways exist to leverage this to decide on a safe and goal-directed plan for the ego-vehicle.
+In this survey, we systematically categorize and review methods that integrate prediction and planning for self-driving vehicles.
+We highlight their capabilities and limitations and discuss prospects for future research.
+I-A Scope
+In this work, we review methods to integrate prediction and planning into an automated driving system.
 Prediction is the task of anticipating intents and future trajectories of observed traffic agents [8], and planning is about finding the best possible trajectory w.r.t.
@@
 Deep Learning (DL)-based methods follow a data-centric approach to tackle these problems [10] and have led to significant improvements in many fields [11].
-Works in prediction and planning increasingly employ DL-based approaches as well [12, 13, 14, 15] and represent an alternative to traditional, rule-based methods [16, 17] and constrained optimization [18].
+Works in prediction and planning increasingly employ DL-based approaches as well [12, 13, 14, 15] and represent an alternative to traditional, rule-based methods [16, 17, 18] and constrained optimization [19].
 In this work, we focus on DL-based methods.
-The architecture of an ADS can take different forms, which are shown in Fig 1.
-Black-box methods employ a single neural network for perception, prediction and planning (c.f.
-Fig 1(a)), so that depending on the architecture, no boundaries between the modules can be drawn [9, 19, 20].
-In contrast, modular systems consist of clearly defined components for different tasks [21, 22, 23, 24, 25].
-Traditionally, they are arranged in a sequential order (c.f.
-Fig 1(b)).
-In this work, we emphasize the shortcomings of this approach and highlight works that integrate prediction and planning in a more sophisticated way (c.f.
-Fig 1(c)) to enable bi-directional interaction of self-driving vehicles and surrounding traffic.
-We refer to the component of an automated driving system that integrates prediction and planning as an Integrated Prediction and Planning System (IPPS).
-We limit the scope of our work to mixed traffic without Vehicle-to-X communication between traffic agents.
-Furthermore, we exclude pedestrian motion forecasting, which poses a different problem that is subject to weaker dynamic constraints compared to vehicles [26] and has been thoroughly surveyed [27, 28, 29].
-1.2 Contribution and Structure
-In a thorough approach, this survey first reviews DL-based prediction and planning before focusing on how to integrate these two tasks.
-While surveys on classical methods [30, 31, 14, 32], stand-alone DL-based prediction [33, 34, 35, 15] or planning [36], and end-to-end (E2E) AD [37] already exist, to the best of our knowledge we are the first to review the integration of DL-based prediction and planning.
-We summarize our contributions as follows:
-  • We provide a comprehensive overview of common representations, architectures and state-of-the-art methods for DL-based prediction (Sec. 3) and planning (Sec. 4).
-  • We propose a categorization for the integration of prediction and planning, which addresses model architecture, system design and behavioral aspects.
-Further, we investigate and analyze the interrelations between these categories and their impact on safety and robustness (Sec.
-5).
-  • We reveal gaps in state-of-the-art research and point out promising directions of future research based on the identified categorization (Sec. 6).
-(a) Black box ADS
-(b) ADS with sequential IPPS
-(c) ADS with bidirectional IPPS
-Figure 1: System architectures of automated driving systems.
-Dashed lines indicate, that due to the black-box nature of some E2E ADS, it can be unclear which network components do prediction and planning, so that no boundary around the IPPS can be drawn.
-2 Task Definitions
+An overview of different architectures of automated driving systems is shown in Fig.
+2.
+It can take three forms ranging from modular systems over end-to-end (E2E) differentiable modular systems to monolithic E2E systems.
+Modular and modular E2E systems consist of clearly defined components for different tasks [20, 21, 22, 23, 24].
+Traditionally, they are arranged sequentially.
+We will discuss these systems in Sec.
+III.
+In contrast, monolithic E2E methods employ a single neural network for perception, prediction, and planning so that depending on the architecture, no boundaries between the modules can be drawn [9, 25, 26].
+As the order of the subtasks cannot be determined, we name this integration of prediction and planning “undirected” (c.f.
+Sec.
+IV).
+In this work, we emphasize the shortcomings of sequential and undirected approaches and highlight works that integrate prediction and planning in a more sophisticated way to enable bidirectional interaction of self-driving vehicles and surrounding traffic (c.f.
+Sec.
+V).
+In the following, we refer to the component of an automated driving system that integrates prediction and planning as an Integrated Prediction and Planning System (IPPS).
+Figure 2: Overview of automated driving systems.
+There are three system architectures shown on the left.
+Modular systems consist of individual modules, whose interfaces provide interpretability but restrict information flow and end-to-end differentiability.
+In contrast, monolithic E2E systems are end-to-end differentiable but not interpretable.
+Modular E2E systems combine both properties: they are end-to-end differentiable and interpretable.
+All three system architectures comprise an integrated prediction and planning system (IPPS), depicted in green.
+Our work focuses on this very part and identifies three paradigms to integrate prediction and planning, as shown on the right.
+Sequential IPPSs condition one task on the other.
+Undirected IPPSs allow for more complex interactions but provide low interpretability.
+Bidirectional IPPSs explicitly ensure that both tasks are mutually conditioned on each other.
+Our review analyzes these integration paradigms and highlights their compatibility with different system architectures.
+I-B Contribution and Outline
+By discussing automated driving systems with a focus on the integration of prediction and planning, we take a novel perspective.
+IPPSs can be classified into three categories as depicted in Fig.
+2.
+Accordingly, this survey is structured into three main chapters: sequential (Sec.
+III), undirected (Sec.
+IV), and bidirectional IPPSs (Sec.
+V).
+While surveys on classical methods [27, 28, 14, 29], stand-alone DL-based prediction [30, 31, 32, 15] or planning [33], and E2E AD [34] already exist, we observe that recent methods thoughtfully design the interplay of prediction and planning [35, 36, 37, 38].
+To substantiate this observation with broad evidence and a theoretical foundation, we summarize our contributions as follows:
+  • We propose a categorization for the integration of prediction and planning based on the dependencies between both tasks (cf.
+Sec.
+IIIIV and V).
+We investigate how these categories relate to system architectures, behavioral aspects, and safety.
+  • In particular, we provide a comprehensive overview of design choices in prediction and planning modules of modular (E2E) architectures (cf.
+Sec.
+III-A and III-B) and discuss their impact on interaction and system-level behavior.
+  • We reveal gaps in state-of-the-art research and point out promising directions of future research based on the identified categorization (Sec. VI).
+II Task Definitions
 In the following, we start by introducing the terminology and notation for the task definitions used in this work.
@@
 $Y_{i}=\{y_{t+1},y_{t+2},\ldots,y_{t+t_{\textrm{pred}}}\}.$ (3)
-and the future states of all SVs are $\overline{\textbf{Y}}_{\textrm{SV}}$. In the following, state sequences are also referred to as trajectories.
+and the future states of all SVs are $\overline{\textbf{Y}}_{\textrm{SV}}$.
+In the following, state sequences are also referred to as trajectories.
 Additional scene information, such as a semantic map or traffic signs and traffic light states are represented by $I$.
-Prediction. Following [8], we state trajectory forecasting as the task of estimating a probability distribution
+Following [8], we state trajectory prediction as the task of estimating a probability distribution
 $P_{\textrm{pred}}=P(\overline{\textbf{Y}}|\overline{\textbf{X}},I)$ (4)
@@
 Some methods omit scene information $I$ and infer trajectories based on a state history alone [43, 44, 45].
-TABLE I: Forms of Prediction
-$\overline{\textbf{Y}}$ $\overline{\textbf{X}}$ Task
-$Y_{\textrm{SV}}$ $\overline{\textbf{X}}_{\textrm{SV}}$, $X_{\textrm{EV}}$ Single-Agent Prediction
-$\overline{\textbf{Y}}_{\textrm{CV}}$ $\overline{\textbf{X}}_{\textrm{SV}}$, $X_{\textrm{EV}}$ Clique Prediction
-$\overline{\textbf{Y}}_{\textrm{SV}}$ $\overline{\textbf{X}}_{\textrm{SV}}$, $X_{\textrm{EV}}$ Joint Prediction
-[…]
-Planning.
-The task is to find a single suitable trajectory for the EV that can be passed on to a downstream motion controller.
+Commonly $\overline{\textbf{X}}$ includes all vehicles in the scene, i.e.
+$X_{\textrm{EV}}$ and $\overline{\textbf{X}}_{\textrm{SV}}$.
+Depending on the vehicles in $\overline{\textbf{Y}}$, different variants of prediction can be formulated: single-agent prediction models the future trajectory for each SV individually [46].
+However, estimating a joint prediction $\overline{\textbf{Y}}_{\textrm{SV}}$ from a set of $n$ single-agent predictions $Y_{\textrm{SV}}$ is not trivial since the number of possible combinations grows exponentially with the number of agents $n$.
+Not all combinations are meaningful, and finding realistic ones with heuristics and joint optimization is cumbersome [47].
+To avoid this, joint prediction directly estimates a joint distribution for multiple SVs [48].
+In planning, the task is to find a single suitable trajectory for the EV that can be passed on to a downstream motion controller.
 Hence, we define planning to be a function $f$ that maps the observational inputs $\overline{\textbf{X}}_{\textrm{SV}}$, and $X_{\textrm{EV}}$ as well as the context information $I$ to a future trajectory $Y_{\textrm{EV}}$, i.e.
@@
 Moreover, it has to be stable in closed-loop, i.e.
-it must be kinematically feasible and result in safe and efficient driving behavior when fed to a downstream controller.
-3 Prediction
-The core of prediction is understanding how the driving scene will evolve.
-Sec.
-3.1 discusses how different scene representations allow local and global interactions.
-Sec.
-3.2 reviews which NN designs are used to model interactions and extract descriptive features.
-Finally, Sec.
-3.3 shows how the extracted features are mapped to trajectory predictions and how multimodality is modeled.
-3.1 Scene Representation
-Figure 2: Comparison of rasterized and sparse representation of the same traffic scene. While the former represents information with fixed-size grids, the latter uses a set of vectors.
-Representing a scene means extracting the relevant subset of all available information and transferring it into a format that subsequent processing steps can utilize.
-In the context of trajectory prediction for automated vehicles, agent states $\overline{\textbf{X}}$ and map I are the most important information to represent.
-Two major representations exist in DL-based approaches: Rasterized and sparse (cf.
+it must be kinematically feasible and result in safe and efficient driving behavior when fed to a downstream controller [49].
+III Sequential IPPS
+To date, the sequential integration of prediction and planning is the prevalent and most widely adopted integration principle.
+Sequential IPPSs execute prediction and planning as separate tasks.
+Prior knowledge is used to design the interplay of the modules.
+Depending on the interfaces, such systems can still be end-to-end differentiable.
+Accordingly, sequential integration can be found in modular and modular E2E systems (c.f.
 Fig.
 2).
-Rasterized.
-Rasterized representations use dense, fixed-resolution grid structures that often have multiple channels [49, 50, 48, 51, 52].
+To investigate different sequential system designs in more depth, we start by taking a closer look at the individual components of prediction and planning and discuss their impact on interaction and system-level behavior.
+III-A Prediction
+III-A1 Input Representations
+In the context of IPPSs for automated vehicles, agent states $\overline{\textbf{X}}$ and map I are the most important information to represent.
+Two major representations exist in DL-based approaches: Rasterized and sparse.
+Rasterized representations use dense, fixed-resolution grid structures that often have multiple channels [50, 51, 52, 53, 54, 55].
 Each channel encodes different information on agent states.
-They are commonly used in combination with rasterized high-definition maps (HD maps).
-Typical modalities are occupancy maps [51], object detections [53], lane markings [54], motion history [55], semantic information [56], traffic signs [57], traffic light states [57], and route information [58] (c.f.
-Fig 2.
-DESIRE [8] is an early work that uses a rasterized representation for DL-based prediction.
-A bird’s eye view (BEV) representation is employed to fuse multiple EV sensor inputs into a shared representation.
-This allows the combination of different sensing modalities and establishes a common coordinate system for all vehicles, facilitating to model interactions within the scene [59].
-Additionally, dense grids are well-suited for processing with powerful Convolutional Neural Network (CNN) [60] architectures [58, 61, 62, 10, 63, 64, 65, 66, 67, 68, 69, 70, 71].
+They are commonly used in combination with rasterized high-definition maps (HD maps) [56, 57] and represented in a bird’s eye view (BEV) [58, 59].
+The BEV allows to fuse multiple EV sensor inputs and different sensing modalities into a shared and interpretable representation [8].
+It further establishes a common coordinate system for all vehicles, facilitating to model interactions within the scene [60].
+Dense grids are well-suited for processing with powerful Convolutional Neural Network (CNN) [61] architectures [56, 62, 63, 10, 64, 65, 66, 67, 68, 69, 70, 71, 72].
 However, embedding all observations in a grid structure leads to information loss due to quantization errors [73].
-Moreover, as discussed in Sec.
-3.2, the locally restricted receptive field of CNNs can hamper interaction modeling.
+Moreover, the locally restricted receptive field and the limited resolution of CNNs can hamper interaction modeling.
 Sparse representations use vectors to describe all objects in a scene.
-The per-object vectors are then either encoded jointly, e.g., with a transformer, or individually and then aggregated, e.g., with Graph Neural Networks (GNN) [73].
-Objects are represented by polygons, or point sets [6, 74, 41, 75, 61, 76, 38, 77].
+The per-object vectors are then either encoded into a latent representation jointly, e.g., with a Transformer, or individually and then aggregated, e.g., with Graph Neural Networks (GNN) [74].
+Objects are represented by polygons, or point sets [6, 75, 42, 76, 62, 77, 39, 78, 79, 80, 40, 81, 82].
 The map is commonly described by a set of lanes represented by polylines and an adjacency matrix.
-Objects and map elements are then encoded into fixed-size latent features, for instance, via Multi-Layer Perceptrons (MLP) [78] or Recurrent Neural Networks (RNN) [79, 39].
+Objects and map elements are then encoded into fixed-size latent features, for instance, via Multi-Layer Perceptrons (MLP) [83] or Recurrent Neural Networks (RNN) [84].
 To encode sequential signals like state histories $\overline{\textbf{X}}$, 1D CNNs [85], RNNs like LSTMs [86] and GRUs [87], and Transformers [88] are applied [42, 89, 77].
-The timeline in Fig.
-3 chronologically depicts impactful prediction methods and exhibits a shift towards sparse representations during the past years.
-Diehl et al.
-[72] proposed a graph representation for vehicle interaction modeling.
-Based on their work, two approaches emerged: Homogeneous and heterogeneous graphs.
-Homogeneous graphs encode objects of one class, for example, the road network [74, 41], whereas heterogeneous graphs combine different object classes in one graph [40, 85, 86, 87].
-However, their work is focused on highway data and does not include map information.
-In contrast, VectorNet [74] proposes to represent all map elements and agents by a set of points.
-Agents are represented by a set of points along their motion history, while lanes are described by polyline points.
-Small objects, such as traffic signs, are represented by a single point, and extended elements are given by polygons defined by multiple points.
-Vectornet then processes each object with a homogeneous subgraph called polyline subgraph.
-A per-object feature is obtained by max-pooling.
-Next, all object features are aggregated in a heterogeneous global graph.
-Finally, the TV trajectory is decoded from the encoding of the node corresponding to the TV.
-Trajectron++ [40] and MFP [88] combine rasterized representations for HD maps with sparse representations for objects.
-Further, a few works use voxel representations that differ in their degree of sparsity [89][90].
-MultiPath++ [38] is the successor of MultiPath [56] and switches from rasterized to sparse representations.
-Their results show that the shift from rasterized to sparse can increase performance.
-The overall development in the field confirms this trend (c.f.
-Fig.
-3).
-Coordinate frame. Irrespective of the scene representation itself, the coordinate frame is a fundamental design choice.
-Applying a global coordinate system with a fixed viewpoint for the whole scene [91, 46, 92, 6] is computationally efficient.
-However, the frame is not viewpoint-invariant, i.e., the predictions change if the coordinate system is moved within the scene.
+Some works like Trajectron++ [41] and MFP [90] combine rasterized HD maps with sparse object representations.
+Throughout the past years, a clear trend towards sparse input representations can be observed in prediction modules [91, 92, 93, 94, 95, 96, 97, 98].
+In the context of integrating prediction and planning, an advantage of sparse representations is that they are object-based.
+Interactions between the EV and SVs can therefore be modeled in a targeted manner.
+This improves the explainability of a system and provides a clear understanding of how surrounding vehicles influence the ego-plan.
+III-A2 Interaction Modelling
+Provided with an accurate scene representation, interaction modeling between perceived objects is a decisive factor for the successful integration of prediction and planning [4, 99].
+Modeling interactions between vehicles helps to understand how the actions of a vehicle affect the behavior of other traffic participants [98].
+This model is subject to various influences, including traffic rules, physics, and common sense reasoning.
+Interaction modeling also includes interactions between map elements and vehicles.
+For example, modeling the relation of vehicles to static map elements like lane markings, traffic signs, etc.
+is important to understand feasible corridors in which trajectories could be localized.
+Some models employ rule-based heuristics to guide interaction modeling, e.g., ScePT [52] only models interaction among nearby vehicles.
+Similarly, heterogeneous graph attention network (HGAT) [100] and graph-based interaction-aware trajectory prediction (GRIP) [101] use heuristics to identify which pairs of nodes in the scene graph are to be connected with edges.
+Other methods rely entirely on neural components to learn these relations.
+Altogether, an interaction model should yield an understanding of how the scene can potentially evolve in the future.
+Recapitulating recent methods exposes that diverse architectures are used to realize interaction modeling.
+RNNs are only used in early prediction models like DESIRE [8] or Trajectron++ [41] and combined with an aggregation operator like spatial pooling or attention.
+Alternatively, CNN methods apply 2D convolutions to implicitly capture interaction within the kernel size [102, 103, 104].
+Compared to sequence processing-based approaches, higher importance is assigned to spatial interaction.
+However, no explicit interactions between designated objects are modeled.
+This shortcoming is addressed by GNNs and Graph-Attention, which explicitly model interactions between individual agents.
+They either combine multiple agents’ features [105] and process them with graph convolution operators [106, 74, 101, 107] or perform graph attention [108, 109, 110, 111, 112] to aggregate information.
+Recently, Transformers [88] are widely adopted for interaction modeling due to their global receptive field and attention mechanism [113, 77, 114, 6, 91, 92, 93, 95, 97, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124].
+All SVs can be predicted simultaneously [77], and the impact of vehicles on each others’ behaviors across different timesteps can be modeled by masking [114].
+Attention is applicable along the spatial and temporal dimension jointly [114], or individually [125].
+The latter break down into sequential [126, 127] and interleaved [6, 92] attention.
+Experiments with Wayformer [92] and SceneTransformer [6] show that joint and interleaved processing outperform sequential approaches.
+As joint processing causes a higher computational burden, the interleaved paradigm balances best between performance and computational demand.
+This property of Transformers can provide explainability to the integration of prediction and planning.
+Attention can be utilized flexibly to achieve locality restrictions, causality, and an explicit interaction design.
+For example, a sequential IPPS could first model interactions between SVs and then include the EV for planning.
+III-A3 Coordinate Frame
+Irrespective of the scene representation itself, the coordinate frame is a fundamental design choice that affects interaction modeling and thus the integration of prediction and planning.
+Applying a global coordinate system with a fixed viewpoint for the whole scene [128, 47, 129, 6] is computationally efficient as it can be shared across both modules.
+However, the frame is not viewpoint-invariant, i.e., the predictions and the plan change if the origin of the coordinate system is moved within the scene.
 This leads to low sample efficiency and impairs generalization [130].
-Even fixing the coordinate origin to the EV center does not lead to viewpoint-invariant predictions when predicting other agents than the EV [94].
-In contrast, other models [95, 96, 74, 54] process a scene from the viewpoint of each agent.
-Thus, the processing is viewpoint-invariant.
-On the downside, computational complexity scales linearly with the number of agents and quadratically with the number of interactions [6].
-An alternative way to achieve viewpoint invariance are pairwise relative coordinate systems introduced in GoRela [97], which describe the relation between agents instead of using a fixed coordinate frame.
+Therefore, the origin is commonly fixed to the center of the EV.
+While this normalization with respect to the EV is a natural choice when planning for the EV, it does not lead to viewpoint-invariant predictions for SVs [131].
+In addition, it complicates the modeling of interactions among SVs.
+For instance, this coordinate frame makes it easy to infer which vehicles are close to the EV or in front of it (small magnitude of $x$ position).
+However, finding SVs that are close to each other is a second-order relation that is harder to learn.
+This is why other models [132, 133, 75, 103] process a scene from the viewpoint of each agent.
+Thus, the processing is viewpoint-invariant and makes it easier to represent interactions between all agents.
+On the downside, these models cannot use a shared representation for prediction and planning, and computational complexity scales linearly with the number of agents [6].
+Moreover, per-agent coordinate frames are unsuitable for joint prediction, as each agent is processed individually.
+An alternative way to achieve viewpoint invariance is the pairwise relative coordinate system introduced in GoRela [94], which describes the relation between agents instead of using a fixed coordinate frame.
 This allows the computation of the relation between static objects offline and greatly reduces computation.
-PEP [94] further reduces computation by using center-relative instead of pairwise relative coordinates, i.e.
+Pioneering Equivariant Planning (PEP) [131] further reduces computation by using center-relative instead of pairwise relative coordinates, i.e.
 shifting the coordinate frame into the mean position of all processed agents before each transformation.
-Another choice to increase robustness are Frenet representations that decompose positions into progress along and lateral offset from a lane [98, 99].
-Either a single Frenet frame defined by the target vehicle is used for all agents [100] or each agent is represented in an individual coordinate frame [84, 101].
-2016201720182019202020212022202320241988DESIRE [8]Fast and Furious $\uparrow$ [102]R2P2 $\downarrow$ [103]MTP [54]IntentNet [49]MATF [89]MultiPath $\uparrow$ [56]MFP [88]CoverNet $\downarrow$ [55]Trajectron++ [40]VectorNet [74]LaneGCN $\uparrow$ [41]TNT $\downarrow$ [75]mmTransformer [76]HOME $\uparrow$ [104]AgentFormer $\downarrow$ [105]DenseTNT $\uparrow$ [106]GOHOME $\downarrow$ [84]MultiPath++ [38]SceneTransformer $\uparrow$ [6]HEAT $\downarrow$ [107]HiVT [108]Wayformer [109]Motion Transformer [110]GoRela [97]LaFormer [111]MotionDiffuser [112]Trajeglish $\uparrow$ [113]BAT $\downarrow$ [114]Scene Representation (3.1)RasterizedSparseMixedInteraction Model (3.2)RNNCNNGNN / AttentionTransformerTrajectory Decoding (3.3)Trajectory RegressionAnchor Trajectory ClassificationEndpoint and CompletionInitial Trajectory and RefineAuto-Regressive FormulationPrediction2016201720182019202020212022202320241988ALVINN [9]Deepdriving [115]End-to-End-CNN [19]FCN-LSTM [116]CIL [117]Camera360 [118]Seg2WP [22]CAL [119]DeepImitative [62]DOCP [10]ChauffeurNet [58]DILP [63]CILRS [64]NMP [61]UD-CIL $\uparrow$ [23]NEAT $\uparrow$ [120]JLTP $\downarrow$ [121]LBC [67]LSD [122]PILOT [123]Transfuser [20]UrbanDriver $\downarrow$ [24]LAV [124]PlanT [125]PDM-Open [126]Input Space (4.1)LatentInterpretableOutput Space (4.2)ActionTrajectoryGoal Conditioning (4.3)NoneSubmodulesInput FeatureRoute CostRoute AttentionPlanning Paradigm (4.4)RegressionCost FunctionHybridPlanning2016201720182019202020212022202320241988PAC [127]HRI [128]PRECOG [65]PiP [129]DSDNet $\uparrow$ [68]P3 $\downarrow$ [66]LookOut $\uparrow$ [69]MP3 $\downarrow$ [70]SafetyNet [130]DIPP [7]SafePathNet [131]GameFormer $\uparrow$ [132]PDM-Hybrid $\uparrow$ [126]UniAD $\downarrow$ [133]TPP [134]VAD $\downarrow$ [135]DTPP [136]Ego-Agents Relation (5.2)UnknownRobot LeaderHuman LeaderJoint PlanningCo-LeaderSafety & Contingency (5.3)No CostMarginalizedWorst-CaseJoint CostContingentGoal Conditioning (4.3)NoneSubmodulesInput FeatureRoute CostRoute AttentionPlanning Paradigm (4.4)RegressionCost FunctionHybridManual IPPS
-Figure 3: Contributions that got attention in the research fields of DL-based prediction, planning, and manual IPPS for automated driving.
-Nodes, stem, and label of each entry encode different aspects that are explained below the timelines.
-Works marked with $\uparrow$ and $\downarrow$ were slightly moved up and down the timeline, respectively, to avoid overlaps.
-3.2 Interaction Modeling
-A decisive aspect of prediction is to model the interaction between traffic agents represented in the scene [4, 137].
-This includes interaction with map elements and other traffic agents.
-Modeling the relation of vehicles to static map elements like lane markings, traffic signs, etc.
-is important to understand feasible corridors in which trajectories could be localized.
-Interaction with other traffic agents are crucial to understanding social interaction.
-RNNs.
-A few early prediction models like DESIRE [8] or Trajectron++ [40] combine RNNs with an aggregation operator like spatial pooling or attention.
-They either extract temporal per-agent features and then aggregate them across all agents or aggregate information first and extract joint temporal features afterwards.
-CNNs.
-Alternatively, Fast and Furious [102], MTP [54], and Multipath [56] applied 2D convolutions to implicitly capture interaction within the kernel size.
-Compared to sequence processing-based approaches, higher importance is assigned to spatial interaction.
-GNNs and Graph-Attention explicitly model interactions between individual agents by combining multiple agents’ features [138] and processing them with graph convolution operators [139, 73, 140, 141] or graph attention [142, 143, 144, 145] to aggregate information.
-Prediction networks like VectorNet [74], LaneGCN [41], TNT [75], HOME [104], DenseTNT [106], GOHOME [84], Multipath++ [38], LaFormer [111], and BAT [114] are examples of this development.
-Transformers [83] have been widely adopted in prediction: InteractionTransformer [146], mmTransformer [76], AgentFormer [105], SceneTransformer [6], HiVT [108], Wayformer [109], Motion Transformer [110], MotionDiffuser [112], Trajeglish [113] and many further works [147, 148, 149, 150, 110, 151, 152, 153, 154, 155, 156] exploit the global receptive field and attention mechanism of Transformers.
-All SVs can be predicted simultaneously [76], and the impact of vehicles on each others’ behaviors across different timesteps can be modeled with attention layers [105].
-Attention layers can be applied along the spatial and temporal dimension jointly [105], or individually [157].
-The latter can be divided into sequential [158, 159] and interleaved [6, 109] architectures.
-Experiments with Wayformer [109] and SceneTransformer [6] show that joint and interleaved processing outperform sequential approaches.
-As, joint processing causes higher computational burden, the interleaved paradigm balances best between performance and computational demand.
-The timeline in Fig.
-3 reflects how interaction models changed together with the scene representation.
-After CNNs and RNNs were applied to rasterized scene representations in early works, GNNs, attention, and Transformers took over since the advent of sparse representations.
-Across all architectures, the prevalent concept is to extend interaction from a local to a global scale.
-HiVT’s ablation study [108] shows that both, local and global attention, are beneficial for prediction and that hierarchical attention yields the best performance.
-3.3 Trajectory Decoding
-The final step of prediction is to generate trajectories from latent features.
-As introduced in Sec.
-2, this step models a probability distribution $P_{\textrm{pred}}=P(\overline{\textbf{Y}}|\overline{\textbf{X}},I)$ over possible future trajectories $\overline{\textbf{Y}}$ given observation $\overline{\textbf{X}}$ and scene context $I$.
-Here, we focus on two major aspects of trajectory decoding: The modeling of multimodality and the decoding principle.
-Multimodality.
-The intention of SVs is inherently uncertain, making $P_{\textrm{pred}}$ multimodal [160].
-Therefore, $P_{\textrm{pred}}$ is commonly either expressed as an explicit density function or estimated with a discrete trajectory set.
-Explicit density functions.
-An intuitive approach to prediction is to explicitly model the probability density function of $P_{\textrm{pred}}$ [61, 66, 69, 70].
-Continuous distributions like bi-variate Gaussian distributions [161], Gaussian Mixture Models [110, 162, 132], or discrete rasterized heatmaps [104, 84] can be found in many prediction works.
+This makes it easier to represent and model interactions.
+It also eases the integration of prediction and planning, as both can use a shared scene representation.
+Thus, they can be combined in a single compute-efficient step.
+The downside of the Cartesian frames mentioned above is that the models are vulnerable to distributional shifts and often produce inadmissible offroad trajectories [134, 135].
+Some works counter this with loss functions that penalize offroad driving [136, 137, 138, 139], driving direction compliance [140] or being distant from the centerline [141].
+An alternative approach is to use Frenet representations that decompose positions into progress along and lateral offset from a lane [142, 143, 89, 144].
+While this improves map compliance (i.e., reduces offroad), it impairs interaction modeling due to the non-euclidean coordinate frame.
+Moreover, this coordinate frame cannot be shared across multiple agents.
+III-A4 Output Representations
+As suggested in Fig.
+2, the interfaces between modules are crucial for the overall system performance.
+To optimize an automated driving system for the final planning task, differentiable interfaces are required [37, 145].
+To provide interpretability, the interfaces should comprise a human-readable representation [62].
+When predicting trajectories in modular E2E systems, the output must fulfill these two criteria.
+Additional difficulty arises from the inherent uncertainty of future prediction [146, 147].
+Since the intentions of SVs are unknown, multiple future modalities must be considered to make a safe plan [5].
+Therefore, the multimodality of $P_{\textrm{pred}}$ is commonly expressed as an explicit density function or estimated with a discrete trajectory set.
+Explicit density functions model the probability density function of $P_{\textrm{pred}}$ [62, 67, 70, 71].
+Continuous distributions like bi-variate Gaussian distributions [148], Gaussian Mixture Models [93, 149, 35, 96], or discrete rasterized heatmaps [150, 89] are used in many prediction modules.
 Alternatively, object-agnostic representations like occupancy maps [151, 152], or flow fields [153, 154, 71] can be used.
-These representations handle multimodality implicitly [61] and were shown to be more robust against perturbations [62].
+Object-agnostic representations handle multimodality implicitly [62] and were shown to be more robust against perturbations [63].
+Further, density functions are interpretable for humans.
 However, since no discrete trajectories are decoded, it is difficult to compare such representations with expert logs for performance evaluation.
-Discrete trajectory sets are either created by sampling from an intermediate distribution or by the model design.
+In an alternative approach to explicit density functions, discrete trajectory sets are either created by sampling from an intermediate distribution or directly output by the model.
 To sample a diverse trajectory set from an intermediate distribution, generative models such as Generative Adversarial Networks (GANs) [155], Conditional Variational Autoencoders (CVAEs) [156], normalizing flows [157] or denoising diffusion [95], can be used [8, 158, 159, 160, 161].
 Alternatively, recent methods apply several prediction heads to the features extracted by a common backbone to decode a set of diverse trajectories [39, 42, 162, 92].
-Other ways to obtain diverse predictions are training loss functions [75], entropy maximization [39, 173], variance-based non-maximum suppression [174], greedy goal sampling [55, 97], divide and conquer strategy [175], evenly spaced goal states [176], or with pre-defined anchor trajectories [56, 177].
+Other ways to obtain diverse predictions are training loss functions [76], entropy maximization [40, 163], variance-based non-maximum suppression [164], greedy goal sampling [165, 94], divide and conquer strategy [166], evenly spaced goal states [167], ensembling [168], or using pre-defined anchor trajectories [104, 169, 170].
 Nevertheless, a guarantee of coverage can only be given by specifying high-level behaviors [171] or regions of the map that must be covered by at least one prediction [77].
-Decoding principles.
-We classify the decoding principles into one-shot, two-shot, and n-shot methods as depicted in Fig.
-4.
-While it is a common practice in planning to decode trajectories as a sequence of actions that translate into Cartesian space using a kinematic model (cf.
+III-B Planning
+III-B1 Input Representations
+The input of the planner depends on the order of modules in sequential IPPSs.
+If prediction is conditioned on potential EV actions, planning is executed first and takes the sparse or rasterized scene representation (cf.
 Sec.
-4.2), only a few works do so in prediction [178, 179].
-Instead, future waypoints are predicted directly from latent features [41, 38, 111, 109, 180].
-Single-shot approaches are split into trajectory regression and anchor trajectory classification.
-Trajectory regression directly decodes the latent features to a trajectory using a NN, typically an MLP [54, 74, 41, 38, 6, 107, 108, 109].
-Especially Transformer-based architectures use direct trajectory regression decoding (cf.
-Fig.
-3).
-However, these methods are prone to predict trajectories that leave the road, cut corners, or are kinematically infeasible [97].
-Specifically designed loss functions for offroad driving [181, 182, 39, 183, 184], driving direction compliance [185] or distance to the centerline [186] can alleviate the problem.
-Anchor trajectory classification states prediction as a classification task.
-Thus, a probability is estimated for a set of pre-defined trajectories [55, 177].
-Defining anchor trajectories can ensure the feasibility of the prediction and allows to impose hard kinematic constraints.
-However, rare trajectories might not be covered by the anchors.
-Two-shot methods structure trajectory decoding into meaningful subtasks that increase explainability and are more flexible than anchor trajectory classification.
-They comprise endpoint and completion as well as initial trajectory and refine strategies.
-Endpoint and completion methods are based on the assumption that a trajectory is mostly defined by its endpoint [25].
-The endpoint can either be regressed [104][84] or classified from a pre-defined set [75][106][97].
-Usually, classification-based methods additionally regress an offset of the classified endpoint [75].
-HOME [104] estimates the endpoint probability distribution as a rasterized map, samples potential endpoints, and then regresses the trajectory using an MLP.
-Initial trajectory and refine approaches combine the ideas of anchor trajectory classification and endpoint and completion.
-Instead of an endpoint, an initial trajectory is classified or generated.
-Then, a refinement step adapts the trajectory, for example, by offset regression for each predicted waypoint [102, 56, 111].
-TPNet [187] and DCMS [188] first regress an endpoint, then generate trajectory proposals and finally refine them.
-N-shot methods are either initial trajectory and refine or autoregressive formulations.
-Initial trajectory and refine strategies in their n-shot variants are similar to the two-shot version.
-The only difference lies in the refinement step, which is realized as a recurrent optimization [8, 112].
-For example, MotionDiffuser [112] samples a noisy trajectory and repeatedly applies denoising diffusion to refine the trajectory.
-Autoregressive approaches use recurrent decoding to perform a stepwise roll-out into the future.
-Iteratively, the next waypoint is predicted, and information on the resulting scene is added to the latent features [103, 189, 40, 76, 105].
-Autoregressive decoding can help to enforce high-quality social interactions and scene understanding [97] but might lead to compounding errors since predicted information is used to make further predictions [190].
-Further, the computational effort and computation time of these methods are usually higher than for one-shot or two-shot methods.
-Fig.
-3 shows the development of trajectory decoding using the categories defined in this Section.
-While endpoint and complete decoding gained popularity, auto-regressive formulations were recently used less often, assumingly due to computation time deficiencies.
-Trajectory regression decoding was always used but became prevalent with the rise of Transformer-based interaction modeling in prediction.
-Figure 4: Principles of trajectory decoding.
-3.4 Benchmarks
-Various prediction datasets provide training, validation, and test splits for benchmarking.
-The most prominent ones are nuScenes [191], WOMD [192], Argoverse [193], Argoverse2 [194], inD [195], highD [196], rounD [197], exiD [198], and Interaction [199].
-Evaluation is carried out by comparing the predicted trajectory of an SV to its ground truth future trajectory.
-To account for the inherent multimodality benchmarks are commonly based on a winner-takes-all evaluation.
-This means the model predicts a fixed number of output trajectories, e.g., six [194] or ten [191], and the evaluation takes only the best one among them into account.
-Alternatively, predictors that output a parametric distribution are often evaluated based on the Negative log likelihood (NLL) of the ground truth under the predicted distribution.
-This assessment also factors in the uncertainty predicted by the model and evaluates the whole distribution instead of only a single trajectory.
-4 Planning
-The planning task is to find a trajectory $Y_{\textrm{EV}}$ for the EV that factors in safety, comfort, kinematic feasibility, and goal-directedness based on the observations $X_{\textrm{EV}}$ and $X_{\textrm{SV}}$ as well as additional context $I$ and optionally $\overline{\textbf{Y}}_{\textrm{SV}}$ (cf.
+III-A1) as input.
+If prediction is executed before planning, the predictions are fed to the planner as an additional input in one of the formats described in Sec.
+III-A4.
+In both cases, goal information for the EV is available in the planning module.
+It can have different forms like lane-level route information [172], high-level commands [173, 174], and sparse goal points [173].
+There are four categories how goal information is incorporated into the IPPS: input features, separate submodules, routing cost, and route attention.
+Simply providing goal information as input feature yields no guarantees for goal compliance but is straightforward and widely adopted [175, 176, 26, 171, 177, 178, 23, 64, 65, 179, 22, 180].
+Separate submodules are used only with high-level commands, which serve as a switch between command-specific submodules [175, 21, 181, 68, 171].
+This can prevent the model from over-adapting to driving modes that dominate the training dataset, but requires a fixed number of high-level commands to be defined in advance.
+A third option to incorporate goal information is by optimizing a cost term, which comprises progress and route compliance [67, 70, 63].
+Balancing multiple targets in a cost term provides flexibility and interpretability but can lead to undesirable tradeoffs with safety.
+Lastly, route attention applies spatial attention to on-route parts of the map.
+This is achieved by Transformer cross-attention [182] or by removing the off-route portion of the map after initial feature aggregation [183, 131, 184, 185, 38].
+Since sequential IPPSs consist of discernible modules, the interfaces between them have to comply with diverse requirements.
+On one hand they should be interpretable to understand failures and facilitate development.
+Rasterized [56, 68, 171, 186] and sparse [187, 188, 184, 189, 23] scene representations (cf.
 Sec.
-2).
-In this section, we will first give a comprehensive overview of common input (Sec.
-4.1) and output representations (Sec.
-4.2).
-Then, we address goal-conditioning in Sec.
-4.3 before we categorize existing works and discuss common paradigms in Sec.
-4.4.
-Finally, we briefly describe existing benchmarks in Sec.
-4.5.
-4.1 Input Representation
-We classify planning input representations into interpretable intermediate representations and latent features.
-Interpretable intermediate representations such as rasterized or sparse scene representations (that were discussed thoroughly in Sec.
-3.1) are common inputs to planning components in modular ADSs.
-Thus, many works exist that build on rasterized [58, 67, 124, 200] or sparse inputs [125, 131, 201, 130, 24].
-Similar to prediction, a clear trend can be observed from rasters to sparse inputs.
-Latent features.
-E2E ADSs directly map sensor information to actions (cf.
-Fig.
-1[9, 202, 20].
-Thus, instead of $X_{\textrm{EV}}$, $\overline{\textbf{X}}_{\textrm{SV}}$, and $I$, the planner’s inputs are latent features.
-The benefit of using learned latent intermediate representations is that no manual engineering effort is needed to design the interfaces between different modules.
-The main drawback of hand-engineered interfaces is that they have to apply to the long-tailed distribution of potential driving scenarios, and every piece of information discarded by the interface cannot be recovered.
-For instance, if a vehicle is currently occluded, but its presence can be inferred from the reflection of its headlight in another car, bounding box representations [58], occupancy fields [70], or affordances [115, 119] will not be able to propagate this information to downstream modules [203, 133].
-In contrast, latent feature representations allow uncertainty propagation, enabling downstream modules to compensate for errors at earlier stages, such as undetected vehicles [61].
-The main disadvantage of latent feature representations is the lack of interpretability.
-In case of failures, evaluating which part of the system caused the error is extremely difficult.
-This aggravates tuning and debugging of pure E2E systems.
-Interpretable E2E systems seek to compensate for this by producing additional intermediate representations [66].
-In contrast to modular stacks, these interpretable intermediate outputs are not used by the planner.
-Instead, they are only used for additional supervision and model introspection.
-The timeline in Fig. 3 reveals that interpretable input representations are becoming increasingly popular compared to latent input features.
-4.2 Output representation
-The planning output is typically represented as a sequence of future states or control actions.
-Future states are given by a sequence of future 2D position and heading angles representing the planned trajectory of the EV over time [22, 62, 58].
-This trajectory is then passed to a downstream controller tasked with tracking it [22].
-The trajectory representation is well interpretable and can be easily checked for collisions, traffic law violations, or deviations from the driveable surface [129, 68].
-Moreover, it is agnostic of the kinematic parameters of the controlled vehicle, enabling generalization to other vehicles.
-However, the ultimate driving performance also depends on the downstream controller and how well the planner and controller are tuned for each other.
-Validating and testing the system can require iterations between both components, which slows down development.
-Future actions are an alternative representation of planned behavior.
-This is mostly adopted by E2E ADSs that regress acceleration and steering angle without the intermediate trajectory representation [117, 118, 119].
-Besides ensuring kinematic feasibility, it can promote comfort as it is directly related to the magnitude of the actions.
-However, the resulting behavior depends on the controlled vehicle’s individual dynamics [204, 19]. Hence, generalization to other vehicles is limited.
-Fig. 3 depicts that early E2E ADSs planned in actions, while trajectory outputs recently gained popularity.
-4.3 Goal Conditioning
-The overall task of ADSs is to navigate safely toward a destination.
-Hence, goal-directedness is one main criterion that determines the suitability of a potential plan.
-Lane-level route information can be provided to the planner by a navigation system.
-Popular benchmarks provide route information through a high-level command such as ’turn left’ or ’go straight’ [205], through sparse goal points along the route [205] or by annotating lanes within the map as on-route or off-route [206].
-For example, in the widely used Carla simulator [205], goal locations are sparsely sampled along the route, and in each planning step, the closest goal location is provided to the planner.
-We define four categories to describe how this goal information is incorporated into the planning algorithm: input features, separate submodules, routing cost, and route attention.
-While the first two were originally introduced in [117], the latter were proposed more recently [201, 207, 126, 208].
-Input features are the most straightforward way to incorporate goal information and have been widely adopted [120, 20, 124, 209, 210, 24, 63, 64, 118, 23, 122].
-Different methods exist to pass this input to the planning module depending on how the route information is represented.
-If the lanes within the map information are annotated as on-route or off-route, then this information can be used as a separate semantic channel in raster maps [58].
-Similarly, vector inputs describing the lane centerlines can have an additional flag reflecting this information [201].
-High-level commands were used as input feature in LBC [124], LAV [67], and AD-MLP [208].
-It was found that inputting high-level intention repeatedly at different stages of the network increases its generalization capability [23].
-Transfuser [20] leverages sparse goal locations as 2D features in the final trajectory decoding step.
-While using goal information as an input feature is simple, goal compliance is not guaranteed.
-On the contrary, [201, 117] demonstrated that the model could ignore this additional input and instead rely on other potentially spurious correlations, such as between the current and the future kinematic state.
-Separate submodules are used only with high-level commands, which serve as a switch between command-specific submodules [117, 70].
-Balancing the dataset to prevent it from being dominated by a single driving mode, such as lane following, is avoided in this setup, as each block is only trained on a single high-level command.
-However, this approach requires a fixed number of high-level commands to be defined in advance.
-Routing cost.
-A third option is to plan trajectories by optimizing a cost term for routing [66, 69, 62].
-Such cost functions can measure the remaining distance to a goal location [62], the progress along or distance to an on-route lane [66].
-Goal-directedness is balanced with further competing objectives such as safety and comfort.
-Thus, the planner can flexibly react to difficult situations, e.g., it could drive off the road to prevent a collision.
-However, tradeoffs between safety and comfort versus progress can be undesirable.
-Route attention applies spatial attention to relevant parts of the map, i.e., the route.
-This can be achieved by removing the off-route portion of the map from the input features, e.g., PDM-Open [126] and PEP [94] discard almost the entire map information and only keep points along a lane centerline on the route.
-Similarly, GC-PGP [201] removes off-route nodes from the lane graph after information aggregation and TPP [134] and DTPP [136] discard all EV trajectory proposals that do not comply with the route.
-Fig.
-3 shows that early models did not condition their plan on a goal at all.
-Afterward, submodules and input features dominated.
-While using input features is easy to implement, it is less effective, whereas submodules represent a major architectural constraint.
-The most interpretable method is using a routing cost function.
-Recent models introduced route attention, which is more flexible.
-Both route attention and routing cost allow to condition on goals that are not reflected in the training dataset, thus improving generalization.
-4.4 Planning Paradigms
-Recall that in Eqn.
-5 we defined planning to be a function $f$ mapping from observational inputs $X_{\textrm{EV}}$, $\overline{\textbf{X}}_{\textrm{SV}}$, $I$ and predictions $\overline{\textbf{Y}}_{\textrm{SV}}$ to a trajectory $Y_{\textrm{EV}}$.
+III-A1) are teherefore common intermediate representations, which are input to the planning module.
+Similar to the input representation used by prediction modules, a clear trend can be observed from rasters to sparse inputs [56, 62, 35, 38].
+On the other hand an interface should convey all relevant information to downstream modules to cover the long-tailed distribution of potential driving scenarios.
+Irrespective of the order of modules, information loss is a problem of hand-crafted interfaces [145].
+For instance, if a vehicle is currently occluded, but its presence can be inferred from the reflection of its headlight in another car, bounding box representations [56], occupancy fields [71], or affordances [190, 181] cannot propagate this information to downstream modules [145, 37].
+Conversely, latent feature representations allow uncertainty propagation, enabling downstream modules to compensate for errors at earlier stages, such as undetected vehicles [62].
+III-B2 Planning Paradigms
+When taking a closer look at the planning module, three general planning paradigms can be observed.
+We introduce each paradigm and discuss its suitability for the integration of prediction and planning.
+Recall that in Eqns.
+5 and 6 we defined planning to be a function $f$ mapping from observational inputs $X_{\textrm{EV}}$, $\overline{\textbf{X}}_{\textrm{SV}}$, $I$ and possibly predictions $\overline{\textbf{Y}}_{\textrm{SV}}$ to a trajectory $Y_{\textrm{EV}}$.
 In the following, we focus on the planning function $f$.
@@
 Hence, in this work, we propose to represent the planning function $f$ as
+$f=h(g(X_{\textrm{EV}},\overline{\textbf{X}}_{\textrm{SV}},I)$ (7)
+or in case of prediction-conditioned planning as
 $f=h(g(X_{\textrm{EV}},\overline{\textbf{X}}_{\textrm{SV}},I,\overline{\textbf{Y}}_{\textrm{SV}})\,.$ (8)
@@
 As planning aims to find a trajectory that optimizes objectives such as safety, comfort, and progress, it is a natural approach to design and optimize a cost function that balances these potentially conflicting objectives.
-Cost function-based planning methods [121, 61, 211] rely entirely on the selection function $h$ to find a suitable trajectory, i.e.,
+Cost function-based planning methods [191, 62, 192] rely entirely on the selection function $h$ to find a suitable trajectory,
 $h=\operatorname*{argmin}_{i}c\left(\hat{Y}^{(i)}_{\textrm{EV}}\right),$ (9)
@@
 Learning-based methods aim to address this by learning a cost function directly from expert demonstrations [201, 202, 203].
-Similarly, [119, 115] learn to regress driving-relevant affordances such as the distance to the leading vehicle or the heading angle relative to the lane, based on which a downstream controller infers driving commands.
 The learned cost function can be non-parametric [62, 204, 205] or comprise hand-designed sub-costs of which only the weights are learned [70, 71, 67, 191].
-The non-parametric cost function of the NMP [61] assigns a cost to each potential future EV position within the planning horizon and is represented as a 3D-BEV-Tensor.
-In contrast, hand-designed sub-costs commonly evaluate safety, comfort, traffic rule compliance, and progress [121, 70, 66].
+Cost function-based planning enables the intuitive integration with a prediction module by including predictions in the cost function [69, 70].
+Depending on the prediction output, this can be realized either by explicit collision checking [189] or by adopting predicted density functions [62] as non-parametric cost.
+For example, the NMP [62] predicts a cost volume, which assigns a cost to each potential future EV position within the planning horizon.
+In contrast, P3 (Perceive, Predict, and Plan) [67] and MP3 (Map, Perceive, Predict and Plan) [71] apply hand-designed sub-costs with learned weighting to heuristically evaluate safety, comfort, traffic rule compliance, and progress of candidate trajectories [191].
+In both cases, the cost function is then used to select the min-cost plan from the candidates.
 Regression approaches rely entirely on the proposal generator $g$.
 The selection function $h$ is the identity, and all burden is placed on the proposal generator $g$, which only generates a single proposal $Y_{\textrm{EV}}=\hat{Y}^{(1)}_{\textrm{EV}}$.
-E2E methods fall into this category.
-Regression approaches can learn to directly predict future waypoints [201, 20, 24] or actions [9] by supervised learning [223].
-This form of imitation learning, i.e., supervised learning from expert data, is called behavior cloning (BC) [224, 190, 225].
-This approach makes the manual design of the cost function obsolete.
-However, BC models are vulnerable to distributional shifts, which occur when the model reaches states not covered by the training data.
-Besides, regressed trajectories are generally not guaranteed to be kinematically feasible.
-UrbanDriver [24] counters this by training the model directly in a closed-loop simulation.
-However, this requires the entire simulator to be differentiable.
-CW-ERM [226] rolls out the policy in a non-differentiable simulation to identify important training scenarios and assign sample weights.
+Typically, regression-based planning is not used in sequential IPPSs but rather found in undirected IPPSs [9, 25, 206, 181, 179, 21, 63, 10, 64, 65, 22, 68, 180, 26, 23, 171, 187].
+Unlike cost function-based planning, regression cannot include predictions explicitly.
+One possibility is to condition the plan on predictions as stated in Eqn.
+8.
+Regression approaches directly output future waypoints  [184, 26, 23] or, more rarely, actions [9, 25, 190, 206].
+Both representations can be converted into one another using a kinematic model [49, 55].
+Commonly, they are trained by supervised learning from expert data [207], called behavior cloning [208, 209, 210].
+This is easy to implement but vulnerable to distributional shifts, which occur when the model reaches states not covered by the training data [23].
+Mitigation strategies include heuristically generated data augmentations or policy rollouts during training [23, 211].
 Hybrid planning describes methods that combine both ideas.
-First, a set of candidate trajectories is regressed by the proposal generator $g$, then the best one is selected with a cost function in $h$ [227].
-For instance, [130, 131, 228, 229, 132] perform collision checks on the proposed trajectories to rule out unsafe proposals.
-In this setting, it often remains unclear which parts of the overall objective are enforced by the proposal generator and which by the cost function.
+First, a set of candidate trajectories is regressed by the proposal generator $g$.
+Then the best one is selected with a cost function in $h$ [212].
+For instance, [189, 188, 213, 214, 35] perform collision checks on the candidate plans to rule out unsafe proposals.
+In contrast to cost function optimization, the generated trajectory proposals are not fixed but scene-dependent.
+III-C Integration
+III-C1 Human Leader
+Figure 3: An exemplary sequential IPPS that falls in the human leader category.
+First, the SV’s future is predicted, and then a plan is made based on the anticipated future.
+In this example, the planning step involves evaluating three candidate plans and selecting the one that does not collide with the predicted SV positions.
+Most sequential IPPSs belong to the human leader integration principle [62, 63, 71, 67, 69, 70, 189, 37, 215].
+Human leader IPPSs can only be realized in modular or modular E2E automated driving systems since the relation of prediction and planning is specifically engineered.
+The planned EV trajectory is conditioned on the predicted SV behavior as depicted exemplarily in Fig.
+3.
+This means the influence of the EV’s plan on the SVs is not modeled.
+Due to this unidirectional dependency, the EV shows a reactive behavior, which can lead to underconfident plans [5].
+In the lane change example in Fig.
+3, the EV tries to find a plan that is suitable given the three predicted SV behaviors without being aware that it can influence them.
+Consequently, it will favor the conservative plan and fail to reach its goal lane.
+If the EV could reason about the influence of its own actions on the predicted SV behavior, it could anticipate that the SV might slow down or speed up further in reaction to its behavior, which would make the lane change maneuver possible.
+This simple example demonstrates the limitations of unidirectional, sequential IPPSs and stresses the advantage of predictions that depend on the EV’s behavior.
+There is a vast body of literature on models that use a human leader sequential IPPSs.
+Deep structured self-driving Network (DSDNet) [69] uses a modular E2E system architecture.
+Each module builds on the outputs of the preceding module and can additionally access the latent feature of the perception backbone.
+Within the IPPS, a set of potential future trajectories is predicted for each detected vehicle.
+The planning module then samples candidate EV trajectories and applies a traditional hand-crafted cost function.
+The cost includes the collision probability based on the SV predictions.
+In contrast, the Neural Motion Planner (NMP) [62] is based on a learned cost function.
+Features of a shared backbone network are fed to a prediction module that forecasts future bounding boxes and a planning module that generates a spatio-temporal cost function.
+As in DSDNet, planning involves sampling feasible trajectories, evaluating them with the cost function, and selecting the best one.
+Beside P3 [67], LookOut [70], and MP3 [71] are further examples of this development.
+Specifically, P3 and MP3 predict an occupancy map while LookOut [70] predicts discrete trajectories.
+Then a cost function that combines rule-based and learned components is used to evaluate pre-defined candidate plans.
+In comparison, SafetyNet [189] follows a unique approach.
+A monolithic E2E planner with implicit prediction (cf.
+Sec.
+IV-A) is applied in parallel to a module that outputs explicit predictions.
+Collision checks between the EV plan and SV predictions are performed to ensure the plan’s safety.
+If the initial plan is classified as unsafe, a hand-crafted fallback layer generates a lane-aligned trajectory.
+Hence, the prediction module affects the selected plan, and we classify SafetyNet as a human leader IPPS.
+UniAD [37] implements a completely learning-based modular E2E system, which employs queries as the interface between the submodules.
+Thus the planning module can attend to agent-level features from previous network layers.
+Various works follow a similar approach: Vectorized Autonomous Driving (VAD) [36] and Hydra-MDP [216] use the output features of previous layers explicitly for collision and map compliance checks during training.
+III-C2 Robot Leader
+Figure 4: An exemplary robot leader IPPS.
+First, three candidate plans are generated, e.g., with a kinematic sampler.
+Then, a prediction conditioned on each candidate plan is inferred.
+In the example, the EV expects the SV to brake for it if it slowly merges to the left lane.
+Thus, it opts for this plan.
+Selecting this plan without hedging against the risk that the EV does not brake as expected results in overconfident behavior.
+In robot leader IPPSs, candidate EV plans are inferred based on the current state, and the prediction step for the whole environment is conditioned on it [217].
+For instance, [218, 219, 217] predict SV trajectories conditioned on candidate plans for the EV with a neural module and then select the best one with a hand-crafted cost function.
+Intuitively, these methods forecast different future scenarios based on pre-defined candidate plans.
+Like human leader IPPSs, they can only be realized in modular or modular E2E automated driving systems due to the manually engineered influence of EV and SVs.
+In contrast to human leader IPPSs, the EV can anticipate the reaction of SVs to its own actions in robot leader IPPSs.
+Hence, the EV can effectively seek to make the SV react to its trajectory, which can result in aggressive driving behavior [220, 221, 90].
+Consequently, in the example in Fig.
+4, the EV might reason that when it follows the fast-progressing plan B, the observed SV will slow down to prevent a collision.
+Overall, robot leader IPPSs are no common choice in AD (cf.
 Fig.
-3 shows that regression is the dominating planning paradigm for E2E systems.
-Cost function-based planning is mainly used in the work of researchers from Uber ATG and the University of Toronto [68, 66, 69, 70].
-4.5 Benchmarks
-Different methods exist to evaluate planning approaches, namely open-loop evaluation and closed-loop simulation.
+7).
+In a pioneering work [220], classic optimization is mixed with learned elements to model the effect of EV behavior on SVs.
+The first fully learning-based method applies a conditional variational autoencoder to obtain diverse rollouts of the future of the scene recurrently [221].
+At each timestep of the rollout, the prediction is conditioned on the next step of a pre-defined candidate plan.
+Evaluating multimodal predictions for each candidate plan with a hand-crafted cost function yields the best plan.
+The example in Fig.
+4 reveals that robot leader systems can, other than human leader systems, anticipate the reaction of SVs to EV actions.
+However, this paradigm might lead to overconfident behavior.
+Altogether, this section has shown that no sequential IPPS can account for the bidirectional influence of EV behavior on SVs and vice versa.
+To resolve the issue, other integration principles are necessary.
+IV Undirected IPPS
+In contrast to sequential IPPSs, which are separated into various submodules, monolithic E2E systems employ a single neural network.
+Implicit systems directly map the state inputs $\overline{\textbf{X}}_{\textrm{SV}}$, $X_{\textrm{EV}}$, and $I$ to the planned EV trajectory $Y_{\textrm{EV}}$ [222].
+Alternatively, joint optimization methods perform prediction and planning in a single step.
+Alongside $Y_{\textrm{EV}}$ they output SV predictions $\overline{\textbf{X}}_{\textrm{SV}}$.
+In the following, we first examine implicit systems and then discuss joint optimization approaches.
+IV-A Implicit
+Reasoning about the future outcome of the environment is crucial when planning the behavior of the EV.
+Hence, it can be assumed that even models without a dedicated prediction output anticipate the behavior of their environment.
+Generally, the expert’s actions, which the E2E planner learns to imitate in training, are based on reasoning of how the future will unfold.
+By learning to imitate actions, the model could exhibit behavior that is implicitly based on such reasoning as well.
+For instance, consider a scenario where the EV’s leading vehicle brakes.
+An expert or a human driver will base their driving decision on the likely future scenario that the safety margin will shrink and thus brake as well.
+Hence, an E2E planner that imitates the expert’s behavior makes a decision that is implicitly based on a predicted future.
+For instance, HydraMDP [216] is a monolithic system that plans by regressing a score for each candidate trajectory within a predefined set.
+In training, an auxiliary collision cost is predicted for each candidate.
+Predicting this accurately is impossible without reasoning about the future behavior of SVs.
+Thus, the SVs’ future behavior and the interactions among them, as well as between the SVs and the EV, are modeled implicitly.
+Pioneered by ALVINN [9], there is a vast body of literature on IPPSs based on monolithic E2E planners [25, 206, 175, 179, 21, 63, 10, 64, 65, 22, 68, 180, 223, 23, 224].
+The main disadvantage is their black-box nature, which makes model introspection and safety verification difficult.
+Moreover, understanding the limitations and their capabilities with respect to their interactive behavior is difficult.
+In fact, this can only be evaluated empirically in comprehensive benchmarks, which is what we outline as a direction for future research in Sec.
+VI-C3.
+To improve interpretability, some works generate interpretable intermediate outputs such as online HD maps [225], semantic segmentations [176], and object detections [26].
+These provide additional training supervision and introspection into the perception part of the system.
+We clearly separate this from systems that additionally output predictions.
+For instance, if the prediction output is branched off from an intermediate feature layer, the downstream part is focused exclusively on planning [56, 62, 176, 187], while the upstream part is relevant to all subtasks.
+Given that this makes planning a downstream task of prediction, the architecture would fall into the human leader category.
+In contrast, if prediction and planning are both outputs of the final layer, we assume that both tasks are addressed in a single step, which means that the architecture resembles a joint optimization system, discussed in the next section.
+IV-B Joint Optimization
+Figure 5: An exemplary joint optimization IPPS.
+A joint global cost function is optimized by selecting the best of several potential scenarios.
+Therefore, a joint prediction and planning step forecasts several potential scenarios that describe the EV and SV behavior.
+Then, each of them is evaluated using a global cost function that evaluates the outcome for all vehicles.
+Here, a plan for the EV is selected that corresponds to the scenario where the SV brakes to let the EV merge before it.
+Joint optimization describes IPPSs that deterministically approximate a joint objective [226] under the assumption that an optimal outcome exists.
+The EV’s plan is obtained by global optimization across all agents [227].
+This erroneously implies that the behavior model used for SVs is exactly known and that every traffic participant optimizes this same global objective.
+This becomes particularly challenging when considering goal-conditioning for the EV.
+If the joint prediction and planning process is conditioned on the EV’s goal, it suggests that all SVs will behave in a manner that enables the EV to achieve its goal.
+At the architecture level, these systems predict and plan in a joint step.
+This can mean explicitly optimizing a learned or hand-crafted cost function or decoding a joint prediction for all agents, including the EV.
+In the latter case, the optimization can be an iterative process [228, 35] or limited to one step [188].
+An exemplary architecture is shown in Fig.
+5, where a global cost function is used to identify the optimal plan for the EV from a set of possible future scenarios.
+The selected plan relies on the SV to break so the EV can merge.
+However, the EV selects this plan without considering that the SV might not behave accordingly.
+In particular, joint optimization applies to monolithic planners that map all inputs to a plan for the EV $Y_{\textrm{EV}}$ and the predicted future of the SVs $\overline{\textbf{X}}_{\textrm{SV}}$.
+As only the EV plan is propagated to a downstream controller and the predictions are not used further, they act as a regularization which can make the learning process more sample-efficient and can improve generalization capabilities [56, 171, 181] as it was previously demonstrated for other auxiliary tasks in prediction and planning like and occupancy forecasting [67].
+A seminal work in this area is GameFormer [35].
+Inspired by hierarchical game-theoretic frameworks [229, 230, 231, 232], it models interaction between agents as a level-$k$ game [233, 234].
+Thus, a Transformer decoder models the interaction between all agents, including the EV, by iteratively updating the individual predictions based on the predicted behavior of all agents from the previous level.
+Similarly, SafePathNet[188] employs a Transformer module for joint prediction and planning of multiple scenarios and then ranks them based on their estimated probability and collision avoidance.
+Generative end-to-end autonomous driving (GenAD) [235] adapts a generative decoder to regress SV and EV trajectories jointly.
+The main problem with this approach is that there is no guarantee that the SVs will behave accordingly.
+Demonstrations show how this can lead to fatal errors [5].
+However, in applications with vehicle-to-vehicle or vehicle-to-infrastructure communication, intentions can be communicated between agents, and global optimization is a valid modeling assumption.
+There, a global cost function reflecting each actor’s individual goals is optimized either by a central server or by negotiation among agents.
+The resulting plan is assigned to each vehicle leading to more efficient and safe traffic [236, 237].
+V Bidirectional IPPS
+Figure 6: Exemplary tree-based co-leader planner.
+Starting at an initial state ($t=0$) a tree is constructed to model potential EV plans and SV behaviors.
+The tree spans across several stages, each reflecting a discrete timestep within the planning horizon.
+In each stage, child nodes are generated by predicting the SVs and unrolling one EV action, both conditioned on the state of the parent node.
+This is visualized in the “rollout step” section.
+Note that the state in the parent node ($t=1$) includes how far the EV has progressed along the candidate plan until this point.
+Thus, the prediction can only observe the candidate plan up to $t=1$.
+It is not conditioned on the EV’s next action, i.e., it’s causal.
+Due to space limitations, we only visualize a prediction with two modes and only a single EV plan per rollout step.
+In practice, a tree might comprise more predictions and evaluate multiple potential EV behaviors.
+Based on the anticipated outcome of each EV behavior, i.e.
+the evaluation of each node w.r.t.
+collisions, progress and comfort, the best action can be chosen by the planner.
+We term systems that react to anticipated behaviors of SVs but also expect SVs to react to their own actions bidirectional IPPSs, as they model the influence of prediction on planning and vice versa.
+Intuitively, this causes a chicken-and-egg problem, where predictions are needed to plan an appropriate behavior, and at the same time, the EV’s future plan is needed to predict the SVs’ reactions to it.
+In this section, we highlight how such systems can be realized by predicting a reactive policy instead of a fixed behavior for SVs.
+The resulting interactive behavior is termed co-leader.
+V-A Co-leader
+Co-leader planning sets the highest requirements for an IPPS.
+It must model the influence of SVs’ potential future behaviors on the EV plan as well as their stochastic reaction to potential ego trajectories [238, 239, 5].
+The result is an interaction-aware system that can model the anticipated behavior of its surroundings and how it can influence it.
+However, it does not assume this reaction to be deterministic, which is in contrast to joint optimization.
+Thus, the EV cannot be certain about SVs’ reactions and must be prepared to react to multiple future outcomes [240].
+For instance, in the opening example in Fig.
+1, the EV can try to approach the intersection to provoke a reaction of the SV.
+This reaction can either indicate that the SV is going to make a turn, e.g., with a turn signal or by slowing down, or the opposite, i.e., accelerating or keeping its speed, to indicate that the SV will continue straight.
+The former will enable the EV to confidently decide in favor of the faster-progressing plan.
+Even though this paradigm is most capable in theory, only a few works exist that fulfill the challenging requirements set to the system.
+V-A1 IPPS Architectures
+Contingencies from Observations (CfO) [5] originally introduced this paradigm, leveraging normalizing flows to represent multi-agent behavior as a learned coupling of independent variables.
+However, the intricate nature of this architecture may pose obstacles to wider adoption.
+Furthermore, their benchmark is simplistic, encompassing only a few scenarios with hard-coded agent behaviors.
+Alternatively, tree-structured methods overcome the circular dependence of prediction and planning by representing future outcomes in a tree, where each node is a state of the EV and its surroundings.
+Multimodality caused by different EV decisions or multiple potential reactions of SVs is handled by branching [241, 242].
+Tree-structured policy planning (TPP) [185] presents a readily comprehensible tree-structured framework that implements co-leader planning.
+It involves sampling kinematically feasible trajectories for the EV and then conditioning prediction on each of them.
+An autoregressive formulation for prediction ensures causality: it can be seen as a way to roll out the scene with SV predictions and the EV plan such that both can only access the observation history up until the respective rollout step.
+Autoregressive decoding can help to enforce high-quality social interactions and scene understanding [203, 243, 41, 77, 114, 94] but might lead to compounding errors since predicted information is used to make further predictions [209].
+Finally, the best EV trajectory proposal is selected with a dynamic programming algorithm.
+Differentiable tree-structured policy planning (DTPP) [38] extends this with a Transformer-based conditional prediction module, which allows the forecast of all agents conditioned on multiple EV plans at once.
+It also replaced the dynamic program with a learnable cost function and applied node pruning to reduce computational burden.
+Still, the EV reaction to the prediction is taken over entirely by the cost function because the sampled ego-proposals are non-reactive with respect to the SV predictions.
+In general, representing the potential EV and SV behaviors in a temporal tree is easy to implement and interpret.
+An exemplary tree-based co-leader planning architecture similar to TPP and DTPP is shown in Fig 6.
+It can be seen that the future planning horizon has to be discretized into a limited number of steps that refer to the levels of the tree.
+Further, the figure shows, how the number of leaf nodes grows exponentially with the branching factor, the depth and the number of candidate plans.
+The branching factor has to be large enough to accurately reflect the uncertainty in the SVs’ behavior as well as the behavior options of the EV and the number of candidate plans can be very high if a dense kinematic sampler is used.
+Moreover, discretizing the future horizon over which a scenario unfolds introduces a sensitive hyperparameter.
+The discretization timestep is subject to a tradeoff between runtime and accuracy w.r.t.
+capturing all relevant intermediate states.
+To counter this, TPP [185] and DTPP [38] build a tree with very limited depth of 2 stages.
+TPP uses a branching factor of 4, while DTPP forecasts 5 states in the first stage and branches each into 6 in the second stage.
+This makes architectures, which are able to model how the future unrolls without relying on discrete timesteps a promising opportunity for future research.
+V-A2 Scenario-Based Trajectory Planning
+Bidirectional IPPSs, i.e., co-leader planning methods, are more complex than sequential or undirected IPPSs because they anticipate and consider multiple future scenarios and to make plans based on this.
+For instance, sequential IPPSs can be used with unimodal predictions or employ a single representation for all scenarios (e.g., occupancy grids), whereas co-leader planning methods must model the multimodal stochastic behavior of their surroundings and plan accordingly.
+Similarly, monolithic E2E IPSSs with undirected integration (cf.
+Sec.
+IV) do not explicitly consider multiple future scenarios.
+Hence, the representation marginalizes all future outcomes.
+While this is easy to realize, it can be safety-critical.
+In the example of Fig.
+1, the cost function must balance high progress and the potential safety risk if the unlikely prediction occurs for one plan with low progress for the other.
+Hence, it must trade off unlikely but dangerous scenarios (e.g.
+collision) against likely low-cost scenarios.
+Conversely, a co-leader architecture must make plans in the face of the uncertain reaction of SVs to its plans.
+In the following, we will discuss methods that incorporate multiple outcome scenarios when choosing among various EV plans.
+We denote the number of relevant future scenarios with $N_{s}$.
+They can be reflected by a multinomial distribution $P\left(\overline{\textbf{Y}}^{(i)}_{\textrm{SV}}\right),i=1,..,N_{s}$ describing the behavior of SVs and the EV in each scenario.
+Based on this, we describe two groups of existing methods: Worst-case planning and contingency planning.
+Their difference lies in the cost function component $h$ of the planning function $f=h(g(X_{\textrm{EV}},\overline{\textbf{X}}_{\textrm{SV}},I)$, which selects one plan among a set of potential plans generated by $g$.
+In the architecture described in Fig 6, $g$ is the tree generation step (green).
+The selection function $h$ would evaluate all nodes and select the most suitable plan (not visualized).
+Worst-case planning refers to IPPSs that evaluate each plan based on the worst-case outcome scenario.
+In the tree-based example, this would mean that if there is a collision in one node, then all nodes that share the same action would be discarded.
+No probabilities $P(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})$ are considered.
+The motivation behind this assumption is that the final output trajectory must be safe in each possible scenario.
+Consequently, the selection function $h$ can be formulated as
+$h=\operatorname*{argmin}_{i}\max_{j}c(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}).$ (10)
+This paradigm strongly focuses on collision avoidance and can result in overly conservative behavior, as shown in Fig.
+1, where only the slower progressing plan is safe under all predictions.
+Thus, it is preferred under worst-case assumptions.
+Contingency planning aims to improve this by additionally taking into account the probabilities $P(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})$ of different future scenarios $\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}$ [244, 245, 246].
+The resulting plan hedges against worst-case risks while enabling progress in expectation.
+For instance, [70, 244, 247] find a plan that is safe for scenarios on the short horizon (independently of their probability, i.e., worst-case) and also optimize the expected cost on the long horizon (marginalization).
+Thus,
+$h=\operatorname*{argmin}_{i}\left(\max_{j}c_{s}(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})+\mathbf{E}_{j}[c_{l}(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})]\right),$ (11)
+where $c_{s}$ and $c_{l}$ consider the short-term and the long-term part of the proposal $Y^{(i)}_{\textrm{EV}}$, respectively.
+In the example in Fig.
+1, this planner allows the most progress while still ensuring safety.
+It can exploit that both candidate plans have a common short-term action.
+Hence, it could opt for the faster-progressing plan as it is aware that it could still re-plan and fall back to the slower one.
+This method can also be applied to non-bidirectional IPPSs.
+For instance, it can be combined with a sequential human leader IPPS, which forecasts multiple scenarios.
+CfO [5] terms this passive contingency, as the EV is contingent upon the SVs [70, 244, 247, 248].
+In contrast, active contingency refers to co-leader systems that are aware that the EV can influence the SVs’ behavior behavior [238, 239, 249].
+VI Discussion and Frontiers
+VI-A Categorization Edge-cases
+In the previous sections, we described and discussed different system architectures and paradigms to integrate prediction and planning in interaction-aware autonomous systems.
+Here, we want to emphasize the distinction between these categories and describe corner cases that transition fluently from one category to another with minor architectural tweaks.
+For instance, an IPPS that directly maps observational inputs to a trajectory for the EV is clearly a monolithic IPPS that falls into the implicit integration category.
+Extending it with an additional decoder head that outputs predicted trajectories for SVs makes it more interpretable.
+For instance, NMP [62] employs such an architecture.
+Thus, the changed architecture consists of a shared backbone for all tasks, which is connected to a prediction decoder module as well as a downstream planner module.
+Therefore, this change transforms the monolithic IPPS into a modular system.
+As the planning step comes after prediction, the system falls into the human leader category.
+Similarly, extending the implicit IPPS architecture with a prediction output from the final layer means that the entire network performs prediction and planning in a joint step, turning it into a joint optimization IPPS.
+Similarly, the exemplary architecture presented in Fig.
+4 can be modified: It is a robot leader IPPS because it expects the SV to react deterministically to the EV’s actions, and the plan is selected with only the EV’s goals in mind.
+However, if the cost function used in the final selection step also optimizes the SV’s outcome, the architecture would fall into the “joint optimization” category.
+VI-B Benchmarking IPPS
+In general, IPPSs need to be evaluated as a system, either because no submodules can be distinguished (monolithic E2E) or because of the interdependence of the submodules.
+VI-B1 Prediction Benchmarks
+A special case are widely adopted sequential human leader architectures, where the prediction component can be evaluated independently of the downstream planning task.
+Various prediction datasets provide expert recordings split into training, validation, and test sets for benchmarking.
+The most prominent ones are nuScenes [250], WOMD [251], Argoverse [252], Argoverse2 [253], inD [254], highD [255], rounD [256], exiD [257], and Interaction [258].
+However, optimizing the prediction performance might not lead to better planning.
+In fact, some works even found that optimizing prediction metrics leads to worse planning results and, thus, worse driving performance [38].
+One reason for this is that prediction metrics average across all target vehicles, i.e., a prediction error for a vehicle far away has the same impact on the score as an error for a vehicle right in front of the EV.
+Moreover, the isolated assessment of prediction modules does not consider the interplay with other submodules.
+For instance, a planner might navigate the EV to states that cause severe distributional shifts and impair prediction performance.
+Thus, it is broadly acknowledged that IPPSs must be evaluated as a system.
+VI-B2 Open-loop Evaluation
+There are two methods to evaluate ADSs as a whole, namely open-loop evaluation and closed-loop simulation.
 Open-loop evaluation is similar to an evaluation in the prediction task.
@@
 More severely, recent work demonstrated that open-loop evaluation results do not correlate with driving performance [183, 261].
-Closed-loop simulation in contrast, lets the planner control the EV.
-Instead of merely comparing the planner output $Y_{\textrm{EV}}$ to that of an expert planner $Y_{\textrm{GT}}$, the evaluation is based on metrics for comfort, safety, and progress.
+VI-B3 Closed-loop Simulation
+Instead of merely comparing the planner output $Y_{\textrm{EV}}$ to that of an expert planner $Y_{\textrm{GT}}$, closed-loop simulations let the planner control the EV.
+The evaluation is based on metrics for comfort, safety, and progress.
 Thus, closed-loop evaluation relates better to real-world driving but requires a simulator to let the model interact with its environment.
@@
 Highway-env is based on simplistic synthetic environments.
+VI-B4 Long-tail Scenarios
 When simulating regular driving scenes, rare and critical scenarios are naturally under-represented.
@@
 Similarly, [5, 65] investigate human-robot interaction more closely with hand-crafted scenarios – an aspect that is uncared for in average driving situations.
+VI-B5 Traffic Agent Simulation
 Another major challenge in closed-loop evaluation is to realistically simulate the SVs.
-Simple simulators such as Highway-Env [262] rely on non-reactive agents that follow predefined trajectories [263].
-In the non-reactive version of nuPlan, and Waymax, these trajectories are obtained from real-world recordings.
+In NAVSIM [288], Waymax, and the non-reactive version of nuPlan, traffic agents follow predefined trajectories obtained from real-world recordings.
 However, if the EV behaves differently than the expert during recording, this can result in collisions caused by the SVs, as they do not react to the EV.
 Hence, this simulation is usually limited to short sequences, where the EV stays close to the expert’s trajectory [289].
-Carla, nuPlan, and Waymax also provide a reactive driver model for the SV, such as the Intelligent Driver Model (IDM) [16, 17, 265, 266, 267, 268].
+Carla, nuPlan, and Waymax provide a reactive driver model for the SV, such as the Intelligent Driver Model (IDM) [16, 290, 291, 292, 293].
+In highway-env, traffic agents can be simulated with a driving model based on IDM+MOBIL [17], which is capable of performing lane changes.
 Another line of work focuses on learning-based methods to simulate realistic driving behavior [294, 295, 296, 297, 298, 299].
 In Waymax [262], users can plug in their own model to control objects in the scene.
-5 Integrating prediction and planning
-Figure 5: Connection between the integration principles, ego-to-agent relationship, and cost function.
-Red arrows for the robot leader planning indicate that applying these cost function categories would be possible but lacks a thorough empirical evaluation.
-We highlight this as an interesting direction for future research.
-After reviewing key aspects such as interfaces and architecture of planning and prediction components found in ADSs, we change perspective and discuss the properties of the IPPS.
-In particular, we will highlight the influence of design decisions in the IPPS on behavior in interactive scenarios.
-We will start by categorizing existing works based on which components the IPPS comprises (Sec.
-5.1).
-Then, we will describe the implications for the interactive behavior which follow from design choices in modular integrated systems (Sec.
-5.2).
-Subsequently, different concepts of safety and contingency in modular integrated systems are discussed in Sec.
-5.3.
-Finally, Sec.
-5.4 discusses possible combinations of these categories.
-An overview is provided in Fig.
-5.
-5.1 Integration Principles
-In this work, we distinguish among three design principles for IPPS. An overview can be seen in Fig. 6.
-Monolithic E2E planning describes methods where the IPPS consists of a single E2E planner, which maps the state inputs $\overline{\textbf{X}}_{\textrm{SV}}$, $X_{\textrm{EV}}$, and $I$ to the EV’s trajectory $Y_{\textrm{EV}}$.
-Thus, the SVs’ future behavior and the interactions among them, as well as between the SVs and the EV, are not modeled explicitly.
-However, the expert’s actions, which the E2E planner learns to imitate in training, are based on reasoning of how the future will unfold.
-By learning to imitate actions, the model could exhibit behavior that is implicitly based on such reasoning as well.
-For instance, consider a scenario where the EV’s leading vehicle brakes.
-An expert or a human driver will base their driving decision on the likely future scenario that the safety margin will shrink and thus brake as well.
-Hence, an E2E planner that imitates the expert’s behavior makes a decision that is implicitly based on a predicted future.
-Pioneered by Pomerlau [9], there is a vast body of literature on IPPS based on monolithic E2E planners [19, 116, 117, 118, 22, 62, 10, 63, 64, 23, 67, 122, 236, 24].
-The main disadvantage is their black-box nature, which makes model introspection and safety verification difficult.
-Interpretable E2E planning frames prediction as an auxiliary learning task [58, 61, 120, 125].
-Explicit predictions are an additional model output trained jointly with the planning task.
-Usually, a shared backbone encodes features that are used for both tasks.
-Then, separate heads decode the respective output representations.
-The additional objective provides a learning signal to the backbone.
-It acts as a regularization which can make the learning process more sample-efficient and can increase generalization capabilities [58, 124, 119] as it was previously demonstrated for other auxiliary tasks in prediction and planning like speed prediction [64], semantic mapping [275], and occupancy forecasting [66].
-However, the additional learning objective results in a trade-off between the different tasks, usually resulting in a hyperparameter that balances the respective losses and has to be tuned empirically.
-Compared to monolithic E2E planning, the explicit prediction adds interpretability and facilitates introspection.
-Nonetheless, both IPPS designs cannot provide safety guarantees.
-Hence, we attribute the increased driving performance to the regularizing effect of the additional objective.
-Manual integration of a planner with a prediction module means that a separate subsystem is employed for each of the tasks.
-The interplay of both tasks is manually engineered based on domain knowledge.
-The most widely adopted integration is a sequential pipeline, where first the prediction module is inferred, and its output is fed to a subsequent planning module [70, 66, 69, 123].
-Since this design is unable to reflect the impact of the EV’s plan on the SVs, an alternative is to swap the order of prediction and planning: First, candidate plans are generated, followed by a conditional prediction module.
-Then, the best plan for the EV is selected among the candidates [129, 134].
-Alternative methods iterate between prediction and planning [276] or do both jointly [132].
-In the following, we give a chronological overview of different models and highlight the integration design of prediction and planning.
-(a) Monolithic E2E IPSS
-(b) Interpretable E2E IPSS
-(c) Manually Integrated IPSS
-Figure 6: IPPS design principles that show how prediction and planning can be integrated.
-Monolithic E2E methods (a) use no explicit prediction, while interpretable E2E methods (b) use it as a regularization in training.
-Manually integrated IPPSs (c) combine a modular prediction with a modular planning method using hand-designed interfaces.
-PRECOG [65] uses a probabilistic model for conditional multi-agent forecasting.
-It implements autoregressive trajectory decoding for EV and SVs jointly.
-At each time step, the state of all agents is updated and used as input for all other agents in the next update step.
-PiP [129, 277] reverses the view of IPPS.
-Its core idea is that SVs behave differently depending on the EV’s actions.
-A trajectory generator provides a set of candidate plans independent of the SVs.
-Each candidate plan is then used to condition prediction.
-Finally, the optimal plan is selected using a cost function.
-DSDNet [68] combines E2E and modular model design.
-Each of the sequential NN modules additionally has access to the high-dimensional features of the perception backbone, complying with the hypothesis [20] that scene context is important for robust planning.
-For each detected vehicle, a set of potential future trajectories is predicted.
-To select the plan for the EV from a sampled candidate set, a hand-crafted cost function is applied, which quantifies the collision probability of the EV candidate trajectories and the SVs’ predicted trajectories.
-P3 [66], LookOut [69], and MP3 [70] use cost functions to link prediction and planning.
-P3 predicts an occupancy map, LookOut [69] employs predicted trajectories to evaluate different pre-defined EV candidate trajectories.
-MP3 [70] extends P3 into a mapless approach by predicting an online map.
-A ’dynamic state map’ comprising detections and prediction is used together with the intended route to evaluate potential trajectories.
-SafetyNet [130] parallelly applies a monolithic E2E planner with implicit prediction and an explicit prediction module.
-Multiple collision checks between EV plan and SVs predictions are performed.
-If the initial plan is classified as unsafe, a hand-crafted fallback layer generates a lane-aligned trajectory.
-SafePathNet[131] improves upon SafetyNet [130] with a transformer module for joint prediction and planning.
-In a mixture-of-experts approach, multiple modes are predicted and ranked.
-The ranking is updated by checking for collisions with the most probable SVs’ predictions.
-Finally, the highest scoring EV mode is selected.
-DIPP [7] jointly predicts the trajectories for all agents, including the EV, which is treated as an SV.
-The mode with the highest probability is selected for each vehicle.
-A differentiable nonlinear motion planner then performs iterative, local optimization under consideration of a kinematic model and a hand-crafted cost function.
-Inspired by hierarchical game-theoretic frameworks [278, 279, 280, 281], GameFormer [132] models interaction between agents as a level-$k$ game [282, 283].
-Thus, a transformer decoder models the interaction between all agents, including the EV, by iteratively updating the individual predictions based on the predicted behavior of all agents from the previous level.
-Hence, the model yields a joint prediction for all agents, including the EV.
-UniAD [133] implements a modular system that is trained E2E based on the planning objective.
-Queries are used as the interface between the submodules so that the planning module can attend to agent-level features from previous network layers.
-Various works adapted or extended this approach: VAD [135] and Hydra-MDP [284] use the output of previous layers explicitly for collision and map compliance checks.
-PPAD [276] repeatedly iterates between prediction and planning based on the perception modules’ outputs and GenAD [285] adapts a generative decoder to regress SV and EV trajectories jointly.
-TPP [134], DTPP [136], and MBAPPE [286] represent multiple future outcomes in scenario trees, where branches reflect combinations of SV predictions and EV plans.
-The best plan is obtained through a tree search.
-Compared to other integration principles, manual integration requires more engineering effort but restricts the solution space in a meaningful way by incorporating prior knowledge.
-Often, manually integrated IPPS provide a high explainability and safer plans than E2E systems [287, 131].
-The timelines in Fig.
-3 show how early works relied mostly on implicit predictions in monolithic E2E designs.
-Starting with ChauffeurNet [58] and PRECOG [65], explicit predictions became more popular.
-More recently, manual integration of E2E differentiable modular components gained popularity [203, 133].
-5.2 Ego-to-Agents Relationship
-After reviewing the integration principles from a system architecture perspective, in the following, we will analyze the manual integration category from a different perspective, namely the intended ego-to-agents relationship.
-This is of special importance in highly interactive scenarios since the EV needs to base its driving decisions on the observed and expected behavior of surrounding agents.
-However, as [5] pointed out, the EV also needs to be aware that it can influence the behavior of others.
-For instance, in the example in Fig.
-7, the plan of the EV can also influence the behavior of the SV: If it approaches the intersection faster, the SV might decelerate and yield to the EV.
-To describe if and in which way the mutual interaction of the EV and surrounding traffic is taken into account in the planning step, [5] introduced the following four categories: robot leader planning, human leader planning, co-leader planning, and joint planning.
-In this context, human refers to the surrounding traffic, and robot refers to the self-driving vehicle.
-In the following, we briefly describe these categories and discuss existing work.
-Robot leader planning.
-The EV plan is inferred based on the current state, and the prediction step for the whole environment is conditioned on it.
-Hence, the EV can effectively seek to make the SV react to its trajectory [88, 128, 127].
-This can result in aggressive driving behavior.
-For instance, in the example in Fig.
-7, the EV would reason that when it follows the fast-progressing plan, the observed SV will yield to it to prevent a collision.
-Human leader planning is the opposite of robot leader planning.
-The plan is based on the forecasted behavior of SVs [61, 62, 70, 66].
-Hence, the EV shows a reactive behavior.
-It does not model the influence of the EV’s plan on the SVs, which can lead to underconfident behavior [5].
-This is the case for IPPS that predict and plan sequentially.
-In the unprotected right turn example (cf.
-Fig.
-7), the EV tries to find a plan that is suitable given the two predicted SV behaviors without being aware that it can influence them.
-Consequently, it will favor the slower plan.
-Joint planning describes systems that are aware that vehicles (including ego) interact with each other.
-The EV’s plan is obtained by global optimization over all agents [288].
-Hence, the IPPS system deterministically approximates a joint objective [289] under the assumption that an optimal outcome exists.
-Hence, it erroneously assumes that the behavior model used for SVs is exactly known and that every traffic participant optimizes this same global objective.
-For example, if the EV in Fig.
-7 squeezes in before the approaching SV, this may be optimal with respect to a reasonable global objective.
-Nonetheless, there is no guarantee that the SV will behave accordingly.
-Consequently, [5] demonstrates how this can lead to fatal errors.
-Co-leader planning is the most complex paradigm.
-It models the influence of SVs’ potential future behaviors on the EV plan as well as their reaction to potential ego trajectories [290, 291, 5].
-In contrast to the joint planning category, the surrounding agents’ behavior is not assumed to be deterministic.
-The EV has to consider this uncertainty in the planning step, e.g., by being prepared for multiple future outcomes and reacting accordingly.
-Additionally, co-leader planning expects the SV to make plans for different potential behaviors of the EV.
-The EV can exploit this by behaving in a way that makes the SVs reveal their intention, which in turn lets the EV make more confident plans.
-For instance, in the example in Fig.
-7, the EV can try to slowly approach the intersection to provoke a reaction of the SV.
-This reaction can either indicate that the SV is going to make a turn, e.g., with a turn signal or by slowing down, or the opposite, i.e., accelerating, to indicate that the SV will continue straight.
-The former will enable the EV to confidently decide in favor of the faster-progressing plan.
-CfO [5] originally introduced this paradigm, leveraging normalizing flows to represent multi-agent behavior as a learned coupling of independent variables.
-However, the intricate nature of this architecture may pose obstacles to wider adoption.
-Furthermore, their benchmark is simplistic, encompassing only a few scenarios with hard-coded agent behaviors.
-TPP [134] presents a readily comprehensible tree-structured framework that implements co-leader planning.
-It involves sampling kinematically feasible trajectories for the EV and then conditioning prediction on each of them.
-An autoregressive formulation for prediction ensures causality: It can be seen as a way to roll out the scene with SV predictions and the EV plan such that both can only access the observation history up until the respective rollout step.
-Finally, the best EV trajectory proposal is selected with a dynamic programming algorithm.
-DTPP [136] extends this with a transformer-based conditional prediction module, which allows the forecast of all agents conditioned on multiple EV plans at once.
-It also replaced the dynamic program with a learnable cost function and applied node pruning to reduce computational burden.
-Still, the EV reaction to the prediction is taken over entirely by the cost function because the sampled ego-proposals are non-reactive with respect to the SV predictions.
-Fig.
-3 shows that no clear trend toward one of the paradigms can be observed.
-While Co-leader planning is theoretically the most capable paradigm, it has not yet been shown empirically if one of these four theoretical concepts is superior to the others and how it can be realized in the design of system architectures.
-This may also be related to a lack of comprehensive empirical benchmarking focused on interactive behavior.
-5.3 Safety and Contingency
-Figure 7: Demonstration of contingency on an unprotected right turn.
-Let’s assume two expected future behaviors of the observed SV, namely $Y^{(1)}_{\textrm{SV}}$ and $Y^{(2)}_{\textrm{SV}}$ with probabilities of $P(Y^{(1)}_{\textrm{SV}})=0.2$ and $P(Y^{(2)}_{\textrm{SV}})=0.8$.
-Two candidate plans for the EV are considered $Y^{(1)}_{\textrm{EV}}$ (purple) and $Y^{(2)}_{\textrm{EV}}$ (orange).
-While the purple plan is only safe if the unlikely prediction $Y^{(1)}_{\textrm{SV}}$ occurs, the orange plan is safe under all predictions.
-Thus, it is preferred under worst-case assumptions.
-When planning with marginalized predictions, the cost function has to balance between the high progress of the purple plan and the potential safety risk if the unlikely prediction $Y^{(1)}_{\textrm{SV}}$ occurs.
-A contingency planner could exploit that both candidate plans have a common short-term action.
-Hence, it could opt for the purple plan as it is aware that it could still re-plan and fall back to the orange one.
-The ego-to-agents relationship categories discussed above highlighted that considering multiple potential future scenarios is essential to safety and contingency planning.
-In the following, we will discuss how this can be incorporated into the cost function component $h$ that we defined as part of the planning function $f=h(g(X_{\textrm{EV}},\overline{\textbf{X}}_{\textrm{SV}},I)$ (cf.
-Sec.
-4.4).
-We will analyze the implications of different design choices on safety and contingency.
-To this end, we introduce the following notation: We denote the number of relevant future scenarios with $N_{s}$.
-These scenarios can be reflected by a multinomial distribution $P\left(\overline{\textbf{Y}}^{(i)}_{\textrm{SV}}\right),i=1,..,N_{s}$ describing the behavior of SVs in each scenario.
-Based on how this distribution is reflected in the design of the IPPS, we form the following three groups of existing methods regarding safety and contingency: Planning with marginalized predictions, worst-case planning, and contingency planning.
-They are briefly outlined in Fig.
+VI-B6 Discussion
+Despite recent advancements in scenario generation and traffic simulation, there are no prominent large-scale benchmarks to evaluate the interactive behavior of IPPSs.
+Real-world datasets commonly comprise a limited number of lane changes or unprotected turns, and synthetic datasets with hand-crafted scenarios don’t scale to the needs of comprehensive benchmarking.
+Moreover, leading benchmarks do not report metrics highlighting the strengths and weaknesses of different IPPS architectures concerning their interactive behavior.
+We emphasize this as an important avenue for future research and hope our work inspires researchers to tackle this problem.
+2016201720182019202020212022202320241988ALVINN [9]Deepdriving [190]PAC [220]End-to-End-CNN [25]FCN-LSTM [206]HRI [221]CIL $\uparrow$ [175]Camera360 [179]Seg2WP $\downarrow$ [21]CAL [181]DeepImitative $\uparrow$ [63]DOCP [10]ChauffeurNet $\downarrow$ [56]DILP $\uparrow$ [64]CILRS [65]PRECOG $\downarrow$ [66]NMP [62]UD-CIL $\uparrow$ [22]JLTP $\downarrow$ [191]MFP $\downarrow$ [90]LBC [68]PiP [217]LSD [180]DSDNet $\uparrow$ [69]P3 $\downarrow$ [67]LookOut $\uparrow$ [70]MP3 $\downarrow$ [71]Transfuser [26]NEAT $\uparrow$ [176]UrbanDriver [23]SafetyNet $\downarrow$ [189]LAV [171]DIPP [7]PlanT $\uparrow$ [187]SafePathNet [188]DiffStack [145]TPP $\uparrow$ [185]GameFormer $\uparrow$ [35]UniAD $\downarrow$ [37]VAD $\downarrow$ [36]PDM-Open $\uparrow$ [183]DTPP [38]GenAD $\uparrow$ [235]VADv2 $\downarrow$ [147]System ArchitectureModularMonolithic E2EModular E2EIPPSImplicitHumanRobotJointCo-LeaderPlanning ParadigmRegressionCost FunctionHybridIntegrated Prediction and Planning Systems
+Figure 7: Important contributions to DL-based planning.
+We highlight how prediction and planning are integrated.
+The nodes and stem of each entry encode different aspects, which are explained below the timeline.
+To avoid overlaps works marked with $\leftarrow$ and $\rightarrow$ were slightly moved into the past and future, respectively.
+VI-C Challenges and Trends
+In the following section, we discuss existing trends and future challenges in the field of IPPSs.
+Based on our overview of deep learning-based IPPSs and their components, we identify three challenges for future research: system design, comprehensive benchmarking, and vehicle-to-X communication.
+After discussing them, we draw a final conclusion.
+VI-C1 Trends
+A chronological overview and categorization of influential works in the field is given in Fig.
 7.
-Planning with marginalized predictions describes IPPS that do not explicitly distinguish multiple future scenarios $N_{s}$.
-This means that the prediction $\overline{\textbf{Y}}_{\textrm{SV}}$ is marginalized over the future outcomes $\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}$.
-This can be done explicitly, i.e., by integrating over the scenario probabilities $P\left(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}\right)$
-$\overline{\textbf{Y}}_{\textrm{SV}}=\sum_{j}^{N_{s}}\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}\cdot P\left(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}\right)$ (9)
-or implicitly, i.e., by representing all possible scenarios in a joint representation instead of differentiating scenarios.
-For instance, IPPS that incorporate the prediction in their cost function $h$ and do not consider separate future outcomes when evaluating a proposal fall into this category [61, 70, 66].
-Hence, their selection function takes the form
-$h=\operatorname*{argmin}_{i}c(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}_{\textrm{SV}}).$ (10)
-Similarly, this is the case for all systems in the monolithic E2E category (cf.
-Sec.
-5.1[201, 20, 24, 9, 19], as they do not explicitly consider multiple future scenarios and hence – if at all – learn marginalized predictions.
-Broadly speaking, this means that the IPPS assumes that all $N_{s}$ scenarios will occur to a certain degree at the same time.
-As can be seen in the example of Fig.
-7, the cost function needs to trade off unlikely but dangerous scenarios (e.g.
-collision) against very likely low-cost scenarios.
-This is safety-critical since planners need to be cautious, especially w.r.t.
-unlikely but dangerous events.
-Worst-case planning refers to IPPS that are aware that multiple future outcomes $\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}$ exist.
-In this category, no probabilities $P(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})$ are considered.
-Instead, each proposal is evaluated based on the worst-case scenario.
-The motivation behind this assumption is that the final output trajectory must be safe in each possible scenario.
-Consequently, the selection function $h$ can be formulated as
-$h=\operatorname*{argmin}_{i}\max_{j}c(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}).$ (11)
-This paradigm strongly focuses on collision avoidance and can result in overly conservative behavior, as shown in Fig.
-7.
-While this assumption is not a popular choice in DL-based prediction and planning, it is broadly applied for rule-based safety layers such as RSS [292] and can be seen as an intermediate step between planning with marginalized predictions and contingency planning.
-Contingency planning prepares for the unknown future development of the scene by taking into account different future scenarios $\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}$ and their probabilities $P(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})$.
-The resulting plan hedges against worst-case risks while enabling progress in expectation.
-Both the cost function-based and the hybrid planning paradigm can have these properties, and the selection function $h$ takes the general form
-$h=f_{c}(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}^{(j)}_{\textrm{SV}},P(\overline{\textbf{Y}}^{(j)}_{\textrm{SV}}))\,.$ (12)
-where $f_{c}$ is a contingent selection function.
-For instance, [69, 287, 293] find a plan that is safe for scenarios on the short horizon (independently of their probability, i.e., worst-case) and also optimize the expected cost on the long horizon (marginalization).
-Thus,
-$h=\operatorname*{argmin}_{i}\left(\max_{j}c_{s}(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})+\mathbf{E}_{j}[c_{l}(Y^{(i)}_{\textrm{EV}},\overline{\textbf{Y}}^{(j)}_{\textrm{SV}})]\right),$ (13)
-where $c_{s}$ and $c_{l}$ consider the short-term and the long-term part of the proposal $Y^{(i)}_{\textrm{EV}}$, respectively.
-In the example in Fig.
-7, this planner allows the most progress while still ensuring safety.
-CfO [5] further differentiates between passive contingency where the EV is contingent upon the SVs [69, 287, 293, 294] and active contingency where the EV also expects the SVs to be contingent upon it [290, 291, 295].
-5.4 Possible Combinations
-In the previous sections, we have described three ways to categorize the integration of prediction and planning methods, namely (1) integration principle, (2) ego-to-agents relationships, and (3) safety and contingency.
-In the following, we discuss possible combinations of these three dimensions.
-Fig.
-5 shows an overview.
-Our key insight is that the categories we characterized describe the differences between IPPS methods on different levels.
-While the integration principle focuses on a high-level system architecture, the ego-to-agent relationship is mostly based on interactive behavior.
-The safety and contingency consideration is founded on the cost function, i.e., a specific design choice for the proposal selection.
-In the following, we highlight associations between these categories.
-In case of monolithic E2E [9, 20, 24, 19, 117, 64, 10, 121, 23, 67, 122, 62, 22, 118, 116, 201] and interpretable E2E [58, 61, 296], no part of the system architecture explicitly reflects an intended interactive behavior.
-Hence, they cannot be assigned to the ego-to-agents relationship categories, and we determine their ego-to-agents relationship as unknown.
-For instance, imagine that a model is trained on demonstrations from an expert who follows the co-leader planning paradigm and employs contingency planning.
-Given that the model excels at imitating the expert, it could be assumed that the model behaves in the same way.
-However, there is no guarantee for this, and it could be equally possible that the model is unable to reason effectively about SVs, resulting in a behavior similar to the robot leader paradigm.
-Similarly, as E2E models do not use hand-crafted cost functions with a pre-defined structure, the safety and contingency categories described in Sec. 5.3 do not apply to them.
-In contrast, inferring the ego-to-agents relationship of modular systems is often possible.
-Thus, we can further divide manually integrated architectures into robot-leader, human-leader, joint planning, and co-leader planning.
-There is a plethora of manual integration methods following the human leader paradigm (cf.
-Fig.
-3).
-By following a traditional sequential scheme of prediction and planning, [70, 66, 69, 120, 68, 130] do not model the influence of the ego plan on SVs.
-This can be combined with any of the three safety and contingency categories.
-Works like DSDNet [68], P3 [66], MP3 [70], UniAD [133], and  [297] leverage marginalized predictions.
-While applying a worst-case cost function is theoretically possible, most works that can distinguish multiple future scenarios aim to make contingency plans [69, 287, 293].
-Similarly, we argue that the robot leader paradigm can be combined with all three cost functions.
-Consider a simple robot leader model that first identifies potential ego plans and then predicts the future behavior of SVs conditioned on each plan.
-The final selection of the EV’s plan can be based on a cost function belonging to each of the three categories.
-However, existing works use specialized cost functions that do not follow the structure we outlined in Sec.
-5.3, e.g.,[129, 134, 128, 88].
-We want to highlight that systematically combining robot-leader architectures with corresponding cost functions could be a promising direction for future research.
-Especially contingent or worst-case cost functions could mitigate the inherent problem associated with robot leader planning, namely relying on SVs to react in potentially unreasonable ways for the interests of the EV.
-As described in Sec.
-5.2, the active contingency resulting from the co-leader planning paradigm implies that the EV makes contingency plans while being aware that SVs are contingent upon it.
-Hence, it naturally employs a cost function that supports contingency plans.
-Finally, the joint planning paradigm relies on optimizing a joint cost function for all traffic participants in the scene.
-Thus, none of the safety and contingency concepts are applicable, and we denote such a cost function as joint cost.
-6 Outlook and Conclusion
-Based on the overview we gave on deep learning-based prediction, planning, and their integration in automated driving systems, we identify two core challenges for future research: System design and comprehensive benchmarking.
-After discussing both, we draw a final conclusion.
-System Design.
-Employing a traditional, strictly sequential system consisting of perception, prediction, planning, and control is still a popular choice.
-Our survey suggests that this approach is not able to meet the requirements set for driving systems.
-Alternative methods integrate prediction and planning in a way that allows conditioning prediction on potential ego plans and vice versa.
-Nonetheless, it remains unclear which exact integration architecture is most effective.
-Especially in the increasingly popular field of interpretable end-to-end systems, it is unclear how prediction and planning should be integrated.
-Comprehensive Benchmarking.
-We discussed different aspects of the integration of prediction and planning.
+Clustering research trends based on this overview yields three main phases of IPPSs.
+Until 2019 mainly monolithic systems with implicit prediction and output regression were used.
+Following works since 2020 parted with regression and instead incorporated prior knowledge in interpretable cost functions.
+Most of them integrated unconditioned predictions into the cost function and thus followed a human leader approach, which can lead to overly cautious behavior.
+Many works have already modularized the monolithic system and are E2E-trainable.
+Notably, almost all current state-of-the-art systems are end-to-end differentiable.
+In a third phase, beginning at the end of 2022, the integration of prediction and planning is getting more complex.
+In addition to sequential systems with human leader integration, undirected systems with joint IPPSs and bidirectional systems with co-leader IPPSs are implemented more frequently.
+Hybrid planning is also used more frequently than before.
+After cost functions were initially used to evaluate fixed candidate plans, newer methods often regress the candidates based on the scene and then apply the cost function.
+VI-C2 System Design
+Our survey suggests that strictly sequential architectures cannot meet the requirements set for driving systems.
+Nonetheless, it remains unclear which integration architecture is most effective and if the theoretical advantages of bidirectional IPPSs can be realized in practice.
+Especially the interactive behavior has seen little attention so far.
+We believe that E2E differentiable bidirectional IPPSs offer great potential for future study.
+VI-C3 Benchmarking in Interactive Scenarios
+We discussed different aspects of various IPPS architectures and design paradigms.
 However, no comprehensive empirical benchmark reproduces and analyzes their strengths and weaknesses.
-Such an overview would help to better understand the effects of different ego-to-agents relationships and safety/contingency paradigms.
+Such an overview would help to better understand the effects of different system architectures.
 This requires simulation in realistic and highly interactive scenarios with realistic driver models for surrounding vehicles and expressive interaction metrics.
+In this regard, data-driven simulation is a fundamental direction for future research.
+While powerful generative models have already been applied to scenario generation, simulating and evaluating the interactive behavior requires appropriate reactive policies to simulate traffic agents.
 Moreover, benchmarks must cover corner case scenarios and test the generalization to distributional shifts.
+VI-C4 Vehicle-to-X Communication
+The systems described in this work leverage predictions to represent the anticipated intentions of other traffic participants.
+With vehicle-to-vehicle or vehicle-to-X communication, these intentions don’t need to be predicted but can be transmitted to surrounding agents.
+Thus, joint optimization methods (see Sec.
+IV-B) that can optimize several vehicles’ goals can improve traffic efficiency and safety.
+However, such systems also need to hedge against erroneous transmissions or even malicious attacks.
+Moreover, they still need to be able to interact with non-communicating human-driven vehicles.
+Consequently, extending IPPS architectures to incorporate this additional information is a promising direction for future research.
+VI-D Conclusion
 In this work, we surveyed and analyzed the integration of prediction and planning methods in automated driving systems based on a comprehensive overview of the individual tasks and respective methods.
 We described, proposed, and analyzed categories to compare integrated prediction and planning works and highlighted implications on safety and behavior.
+Based on this, we discussed the shortcomings of undirected and sequential IPPSs and highlighted the need for IPPSs that model bidirectional interaction.
+We hope that this way of approaching the architectures of IPPSs opens a new perspective to understanding their implications and capabilities with respect to interaction with traffic.
 Finally, we pointed out promising directions for future research based on the identified gaps.